### v2

# Collaborative Agents for Software Engineering

Abstract
Code review is a heavily collaborative process, which aims at ensuring the overall quality and reliability of software. While it provides massive benefits, the implementation of code review in an organization faces several challenges that make its automation appealing. Automated code review tools have been around for a while and are now improving thanks to the adoption of novel AI models, which help can learn about standard practices and systematically check that the reviewed code adheres to them. Unfortunately, existing methods fall short: they often target a single input-output generative model, which cannot simulate the collaboration interactions in code review to account for various perspectives; they are also sub-performing on various critical code review sub-tasks. In this paper, we advance the state of the art in code review automation by introducing CodeAgent, a novel multi-agent-based system for code review. Fundamentally, CodeAgent is steered by QA-Checker (short for “Question-Answer Checking"), a supervision agent, designed specifically to ensure that all agents’ contributions remain relevant to the initial review question. CodeAgent is autonomous, multi-agent, and Large language model-driven. To demonstrate the effectiveness of CodeAgent, we performed experiments to assess its capabilities in various tasks including 1) detection of inconsistencies between code changes and commit messages, 2) detection of vulnerability introduction by commits, and 3) validation of adherence to code style. Our website is accessed in https://code-agent-new.vercel.app/index.html.
1 Introduction
Code review, as a fundamental activity in software engineering, has been widely studied in the literature [2; 3; 8]. It involves multiple team members, with different expertise and experience, collaborating to check the code along several dimensions, including whether the new code is aligned with existing code in terms of style [14], whether the pull request contents are consistent [43], whether some vulnerability is being injected [4], etc. This review effort is critical for ensuring the stability, quality and readability of the code. Historically, code review has been an intensive manual exercise. Nowadays, however, software projects evolve at a rapid pace with a high frequency of commits and pull requests. For example, in 2022, GitHub, a global platform for software development, witnessed over 3.5 billion contributions, including commits, pull requests, and other forms of contributions [16; 30]. Reviewing these contributions before they are integrated into any code base requires automated tools to ease the workload of code contributors and code reviewers.
In recent literature, various approaches [35; 32; 34] have been proposed to elevate code review automation, using Transformers [36] for code modeling, code abstraction tools such as src2abs [33] and code transformation engines such as AutoTransform [32]. These state-of-the-art approaches however mainly focused on how to rewrite and adapt the submitted code. Code review, however, is actually a collaborative and iterative process involving multiple stakeholders exchanging on a specific thought process [2] and addressing a wide variety of subtasks such as analyzing the consistency between a code change and the associated commit message () [43], detecting potential vulnerability injection () [4], assessing the consistency of code format () [14], etc.
To address the challenges of complex software collaboration eco-system, researchers have applied agent-based systems across a spectrum of tasks, advancing past the conventional single input-output paradigm [40; 38]. Notably, the advent of multi-agent collaboration stands out as a key innovation, especially in simulating human-like behaviors [10; 24; 27] and harnessing the collective strengths of various agents [7; 22; 17].
In this paper, drawing on the success of agent-based collaboration, we developed an agent-based framework, CodeAgent, to simulate the dynamics of a collaborative team engaged in code review processes, incorporating various roles such as code authors, reviewers, and decision-makers. Nonetheless, a notable hurdle within multi-agent systems and Chain-of-Thought (CoT) reasoning is the propensity for conversation topics to drift off course, underlining the necessity for strategies to maintain topic relevance and coherence [20; 6]. The occurrence of drift, often triggered by the model-inspired tangents or the randomness of Large Language Models (LLMs), necessitating the integration of a QA-Checker. This QA-Checker, serving as a crucial supervisory agent within CodeAgent, meticulously monitors the flow of conversation, ensuring that questions and responses remain pertinent and on track, thus maintaining the dialogue’s intended direction [20; 6]. As an instruction-driven entity, the QA-Checker not only refines queries but also realigns answers to match the original intent, employing a systematic approach grounded in a mathematical framework. This involves leveraging a quality assessment function and the Newton-Raphson optimization method [41] to iteratively guide the conversation towards optimal coherence and relevance.
To evaluate the performance of CodeAgent, we employ two distinct sources of data. The first encompasses pre-existing datasets: Trans-Review, AutoTransform, and T5-Review, which have been referenced by leading research in the field. By conducting experiments with these datasets, we aim to compare the code revision capabilities of CodeAgent against those identified as state-of-the-art. The second data source comprises a collection of Pull Requests from GitHub, featuring an extensive array of commits, messages, and comments across nine programming languages post-April 2023. These languages include Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. This diverse dataset is specifically curated for testing format analysis. The experimental results indicate that CodeAgent outperforms the state-of-the-art results in terms of qualitative metrics for code review, achieving a 41.54 percentage point (pp) increase in hit rate (Rate in Table 3) confirming vulnerable issues within merged commits across the nine programming languages. CodeAgent also excels in consistency detection, vulnerability identification, and format analysis, outperforming ChatGPT in overall metrics.
In conclusion, we summarize our contributions as follows:
We build a code review dataset with more than 3,545 real-world commits, commit messages, and corresponding original files for assessing code review model performance in various areas including consistency detection between commits and commit messages, vulnerability detection, code style detection, and code revision.
To the best of our knowledge, we are the first to propose an autonomous agent-based system for practical code review in the field of software maintenance.
Experimental results show that CodeAgent confirms 104 (23.20% higher) more vulnerable issues. Compared to state-of-the-art, CodeAgent improves 5.62pp and 4.00pp of recall and F1-Score, respectively, for the consistency detection between commit and commit message. CodeAgent also boosts 15.96pp and 10.45pp of the recall score and F1-Score, respectively, for the format consistency detection. On the code revision task, CodeAgent surpasses the state-of-the-art from 29.80pp to 31.60pp of the Edit Progress (EP) metric [44] on average.
2 Tasks and Definition
In this section, we summarize our tasks (❶,❷,❸,and ❹) and definitions (,,) in Table. 1.
| Index | Tasks |
| ❶ | Semantic consistency detection between commit and commit message |
| ❷ | Vulnerability analysis |
| ❸ | Format consistency detection |
| ❹ | Code revision |
| Short Term | Definition |
| Consistency analysis between commit and commit message | |
| Vulnerability analysis | |
| Format consistency analysis between commit and original files |
3 CodeAgent
This section details the methodology behind our innovative CodeAgent framework. We first describe the defined role cards in Section 3.1 and discuss the pipeline in Section 3.2. Finally, we will discuss the design of the QA-Checker.
3.1 Role Card Definition
As shown in Figure 1, we defined six characters in our simulation system (CodeAgent), including , , , , , , and they are defined for different specific tasks.
For each, we defined a role card, which contains: 1) The role name is put on the left-upper corner of each card; 2) The phases of the role involved are put on the right-upper corner of each card; 3) On each role card, we show the role-involved conversation and collaborative roles; 4) We show the intermediate output of the role on the right-hand side of the card; and 5) Finally, we put the corresponding files or content out of conversations on the bottom of the card.
All tasks are processed by the collaborative work of two agents in their multi-round conversations. For example, as a role , her responsibility is to do the code review for given codes and files in three aspects (task ❶, ❷, and ❸ in Table 1) and provide a detailed description of observation. ’s code review activity is under the assistance with as shown in Figure 2. Meanwhile, with the Review’s assistance, can process the code revision as shown in the ‘Revised codes’ part in the card in Figure 1. Apart from , also cooperates with and in the simulated team.
3.2 Pipeline
We specifically consider the following sub-tasks of code review: consistency detection between commit and commit message () [43], vulnerability injection detection () [4], Format consistency detection () [14], and code revision [44]. We defined six characters and four phases for the framework. The roles of characters are demonstrated in Figure 1. Each phase contains multiple conversations and each conversation happens between agents. The four phases consist of ① Basic Info Sync, containing the roles of chief executive officer (), chief technology officer () and Coder to conduct modality and language analysis; ② Code Review, leveraging the and for actual code review (i.e., target sub-tasks); ③ Code Alignment, supporting the and to correct the commit by code revision or suggestion feedback to the author; ④ Document, finalizing by synthesizing the opinions of the , , and to give out the final comments.
Apart from six defined roles, the proposed architecture of CodeAgent consists of phase-level and conversation-level components. At the phase level, the waterfall model is used to break down the code review process into four sequential phases. At the conversation level, each phase is further divided into atomic conversations. These atomic conversations involve task-oriented role-playing between two agents, promoting collaborative communication. One agent works as the instructor and the other works as the assistant. Communication follows an instruction-following style, where agents interact to accomplish a specific subtask within each conversation and each conversation is under the supervision of QA-Checker. QA-Checker is used to align the consistency of questions and answers between the instructor and assistant in a conversation to avoid digression. QA-Checker will be introduced in Section 3.3.
Here, we take an example to show the pipeline of CodeAgent. As shown in Figure 2, CodeAgent receives the request to do the code review with the submitted commit, commit message, and original files. In the first phase, , , and will cooperate to recognize the modality of the input (e.g., document, code) and language (e.g., Python, Java, and Go). In the second phase, with the help of , will write an analysis report about consistency analysis, vulnerability analysis, format analysis, and suggestions for code revision. Then, in the third phase, according to analysis reports, will align or revise the code if it finds incorrect snippets with the help of . cooperates with and to summarize the document and codes about the whole code review in the final phase.
3.3 Self-Improving CoT with QA Checker
QA-Checker is an instruct-driven agent, designed to fine-tune the question inside a conversation to drive the generated answer related to the question. As shown in Figure 3, the initial question (task instruction) is represented as , and the first answer of the conversation between and is represented as . If QA-Checker identifies that is inappropriate for , it generates additional instructions attached to the original question (task instruction) and combines them to ask agents to further generate a different answer. The combination in Figure 3 is defined as , where is the additional instruction attached. The conversation between two agents is held until the generated answer is judged as appropriate by QA-Checker, it reaches the maximum dialogue times, otherwise.
Theoretical Analysis of QA-Checker in Dialogue Refinement
The QA-Checker is an instruction-driven agent, crucial in refining questions and answers within a conversation to ensure relevance and precision. Its operation can be understood through the following lemma and proof.
Lemma 3.1.
Let denote the quality assessment function of the QA-Checker for question-answer pair in a conversation at the -th iteration. Assume is twice differentiable and its Hessian matrix is positive definite. If the QA-Checker modifies the question to by attaching an additional instruction , and this leads to a refined answer , then the sequence converges to an optimal question-answer pair , under specific regularity conditions.
Proof.
The QA-Checker refines the question and answers using the rule:
where is the learning rate. To analyze convergence, we consider the Taylor expansion of around :
Substituting the update rule and rearranging, we get:
For sufficiently small , this model suggests an increase in , implying convergence to an optimal question-answer pair as . The convergence relies on the positive definiteness of and the appropriate choice of , ensuring each iteration moves towards an improved quality of the question-answer pair. ∎
In practical terms, this lemma and its proof underpin the QA-Checker’s ability to refine answers iteratively. The QA-Checker assesses the quality of each answer concerning the posed question, employing advanced optimization techniques that is modeled by the modified Newton-Raphson method to enhance answer quality. This framework ensures that, with each iteration, the system moves closer to the optimal answer, leveraging both first and second-order derivatives for efficient and effective learning.
4 Experimental Design
We evaluate the performance of CodeAgent in various qualitative and quantitative experiments in nine programming languages, on four metrics. In this Section, we will discuss experimental settings, including datasets, metrics, and baselines. For more information, please see Appendix B.
4.1 Datasets
For our research, we leverage datasets from prior studies as referenced in the state-of-the-art [44]. Specifically, we utilized Trans-Review, AutoTransform, and T5-Review to assess the edit progress (EP) metric which is also designed by them [44].
As shown in Table 2, in terms of new data collection, which is called codeData, is collected using the GitHub REST API, encompasses over 3,545 commits and 2,933 pull requests from more than 180 projects in nine programming languages (Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby). It focuses on consistency and format detection, featuring both positive and negative samples segmented by the merged and closed status of pull requests across various languages. The detailed information about the dataset can be seen in Appendix-Section E.
| Samples | CA | FA | ||
| Merged | Closed | Merged | Closed | |
| Positive (consistency) | 2,089 | 820 | 2,238 | 861 |
| Negative (inconsistency) | 501 | 135 | 352 | 94 |
4.2 Metrics
F1-Score and Recall. We utilized the F1-Score and recall to evaluate our method’s effectiveness. The F1-Score, a balance between precision and recall, is crucial for distinguishing between false positives and negatives. Recall measures the proportion of actual positives correctly identified [19].
Edit Progress (EP). EP evaluates the improvement in code transitioning from erroneous to correct by measuring the reduction in edit distance between the original code and the prediction. A higher EP indicates better efficiency in code generation [9; 11; 44].
Hit Rate (Rate) We also use hit rate to evaluate the rate of confirmed vulnerable issues out of the found issues by approaches.
4.3 State-of-the-Art Tools and Models
Our study evaluates various tools and models for code revision and modeling. Trans-Review [35] employs src2abs for code abstraction, effectively reducing vocabulary size. AutoTransform [32] uses Byte-Pair Encoding for efficient vocabulary management in pre-review code revisions. T5-Review [34] leverages the T5 architecture, emphasizing improvement in code review through pre-training on code and text data. In handling both natural and programming languages, CodeBERT [12] adopts a bimodal approach, while GraphCodeBERT [13] incorporates code structure into its modeling. CodeT5 [37], based on the T5 framework, is optimized for identifier type awareness, aiding in generation-based tasks. Additionally, we compare these tools with ChatGPT [26] by OpenAI, notable for its human-like text generation capabilities in natural language processing.
5 Experimental Result Analysis
In this Section, we discuss the performance of CodeAgent on four main experiments: vulnerability analysis (Section 5.1), inconsistency detection between commit and commit message (Section 5.2), format inconsistency detection (Section 5.2), and code revision (Section 5.3). We also discuss the difference in execution time of CodeAgent across different languages and conduct capabilities analysis between CodeAgent and other communicative agents in Appendix-Section D.
5.1 Vulnerability Analysis
As shown in Table 1, vulnerability analysis ❷ is one subtask of code reviews. Compared to ❶ and ❸, ❷ is a more complex code review subtask, covering more than 25 different aspects (please see the Appendix-Section F), including buffer overflows, sensitive data exposure, configuration errors, data leakage, etc. This domain necessitates deep technical expertise for accurate data annotation, thereby significantly increasing the time and money cost when labeling manually. Considering the low proportion of commits with vulnerabilities, this paper proposes a proactive verification method for data annotation.
Our approach is twofold. First, we utilize CodeAgent to process 3,545 pairs of commits, commit messages and original files across nine languages. These identified data points for potential vulnerabilities were then subjected to manual verification. Second, we employed tools like CodeBERT [12] and ChatGPT to do vulnerability binary detection in the same dataset. The results were then verified for their authenticity.
| Approach | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | CodeAgent |
| Find | 1,063 | 864 | 671 | 483 | 564 |
| Confirm | 212 | 317 | 345 | 449 | 413 |
| Rate | 19.94% | 36.69% | 51.42% | 92.96% | 73.23% |
| Rate | 5.98% | 8.94% | 9.73% | 12.67% | 11.65% |
Comparison
As delineated in Table 3, the deployment of CodeAgent successfully identified 483 potential vulnerabilities within a dataset of 3,545 items, with an impressive 449 of these ultimately confirmed as high-risk vulnerabilities, substantiated through a rigorous manual verification process exceeding 120 working hours111The verification process involved meticulous manual examination, extending beyond 120 working hours.. CodeBERT, a key pre-trained model for code-related tasks, with its parameters frozen for this experiment, initially identified 1,063 items as vulnerable, yet only 212 passed the stringent verification criteria. Similar trends were observed with ChatGPT-3.5 and ChatGPT-4.0, which confirmed 317 and 345 vulnerabilities out of 864 and 671 identified items, respectively. These outcomes are further quantified by the confirmation rates (Rate) of 19.94% for CodeBERT, 36.69% for ChatGPT-3.5, and 51.42% for ChatGPT-4.0, while CodeAgent demonstrated a remarkable Rate of 92.96%. Additionally, the analysis of confirmed vulnerabilities against all analyzed items (Rate) yielded 5.98%, 8.94%, 9.73%, and 12.67% for CodeBERT, ChatGPT-3.5, ChatGPT-4.0, and CodeAgent, respectively. Evidently, Table 3 not only highlights CodeAgent’s high precision in identifying vulnerable commits but also reveals the progressive improvement from ChatGPT-3.5 to ChatGPT-4.0, likely due to the latter’s capacity to handle longer input sequences, with token limits of 4,096 and 32,768, respectively. The integration of sophisticated algorithms like CoT and QA-Checker in CodeAgent has significantly enhanced its capabilities in vulnerability detection, surpassing the individual input-output efficiencies of ChatGPT and CodeBERT. Further details regarding the importance of the QA-checker can be seen in Appendix-Section C and Section L. Moreover, more experimental results in 9 languages are accessible in Appendix-Section I.
In addition, the analysis of vulnerabilities identified by various models reveals interesting overlaps in their findings. CodeBERT confirmed 212 vulnerabilities, whereas ChatGPT-3.5, ChatGPT-4.0, and CodeAgent confirmed 317, 345, and 449 vulnerabilities, respectively. Notably, the intersection of vulnerabilities confirmed by CodeBERT and ChatGPT-3.5 is 169, indicating a substantial overlap in their findings. Similarly, the intersection between CodeBERT and ChatGPT-4.0 is 170, while a more significant overlap of 212 vulnerabilities is observed between ChatGPT-3.5 and ChatGPT-4.0. The combined intersection among CodeBERT, ChatGPT-3.5, and ChatGPT-4.0 is 137, underscoring the commonalities in vulnerabilities detected across these models. Furthermore, the intersections of vulnerabilities confirmed by CodeBERT, ChatGPT-3.5, and ChatGPT-4.0 with CodeAgent are 212, 317, and 334, respectively, highlighting the comprehensive coverage and detection capabilities of CodeAgent.
Ablation Study
As shown in Table 3, without QA-Checker, CodeAgent is less effective in finding vulnerable issues and reduces the hit rate (Rate and Rate) compared with the full version, indicating the importance of QA-Checker in our CodeAgent. More detailed information about the ablation study can be found in Appendix-Section L.
Moreover, CodeAgent versions 3.5 and 4.0 yielded consistent detection results, but they differ in the quality of explanation when doing the vulnerability analysis of the commit. More details about the difference between CodeAgent-3.5 and CodeAgent-4.0 are provided in Appendix-Section K.2.
5.2 Consistency and Format Detection
In this Section, we will discuss the performance of CodeAgent and baselines on metrics like the F1-Score and recall score of task ❶ and ❸. For ❶ and ❸, the dataset we have is shown in Table 2 and more detailed data information is shown in Figure 7 in Appendix.
Consistency Detection Between Commit and Commit Message
Our comprehensive study, as illustrated in Table 4, assesses CodeAgent’s efficacy in detecting the consistency between commit and commit message, contrasting its performance with other prevalent methods like CodeBERT, ChatGPT-3.5, and ChatGPT-4.0. This evaluation specifically concentrates on merged and closed commits across nine languages, with a keen focus on crucial metrics such as Recall and F1-Score. Notably, CodeAgent exhibits a remarkable performance, outstripping other methods in both merged and closed scenarios. In terms of Recall, CodeAgent achieved an impressive 90.11% for merged commits and 87.15% for closed ones, marking a considerable average improvement of 5.62pp over the other models. Similarly, the F1-Score of CodeAgent stands at 93.89% for merged and 92.40% for closed commits, surpassing its counterparts with an average improvement of 4.00pp. More comparable details in different languages are shown in Appendix-Section. J.
| Merged | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 63.64 | 80.08 | 84.27 | 90.11 | 5.84 |
| F1 | 75.00 | 87.20 | 90.12 | 93.89 | 3.77 |
| Closed | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 64.80 | 79.05 | 81.75 | 87.15 | 5.40 |
| F1 | 77.20 | 87.35 | 89.10 | 92.40 | 3.30 |
| Average | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 64.22 | 79.57 | 83.01 | 88.63 | 5.62 |
| F1 | 76.01 | 87.28 | 89.61 | 93.16 | 4.00 |
Format Consistency Detection Between Commit and Original File
In our detailed evaluation of format consistency between commits and original files, CodeAgent’s performance was benchmarked against established models like CodeBERT and ChatGPT variants across nine different languages. This comparative analysis, presented in Table 5, was centered around pivotal metrics such as Recall and F1-Score. CodeAgent demonstrated a significant edge over the state-of-the-art, particularly in the merged category, with an impressive Recall of 89.34% and an F1-Score of 94.01%. These figures represent an average improvement of 10.81pp in Recall and 6.94pp in F1-Score over other models. In the closed category, CodeAgent continued to outperform, achieving a Recall of 89.57% and an F1-Score of 94.13%, surpassing its counterparts with an average improvement of 15.96pp in Recall and 9.94pp in F1-Score. The overall average performance of CodeAgent further accentuates its superiority, with a Recall of 89.46% and an F1-Score of 94.07%, marking an average improvement of 15.96pp in Recall and 10.45pp in F1-Score. These results underscore CodeAgent’s exceptional capability in accurately detecting format consistency between commits and their original files.
| Merged | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 60.59 | 60.72 | 78.53 | 89.34 | 10.81 |
| F1 | 74.14 | 74.88 | 87.07 | 94.01 | 6.94 |
| Closed | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 69.95 | 73.61 | 68.46 | 89.57 | 15.96 |
| F1 | 80.49 | 84.19 | 80.16 | 94.13 | 9.94 |
| Average | CodeBERT | ChatGPT-3.5 | ChatGPT-4.0 | CodeAgent | Imp (pp) |
| Recall | 65.27 | 67.17 | 73.50 | 89.46 | 15.96 |
| F1 | 77.32 | 79.54 | 83.62 | 94.07 | 10.45 |
5.3 Code Revision
In this Section, we evaluate the effectiveness of various approaches in bug fixing by comparing their Error Percentage (EP) performance. The methods under consideration include Trans-Review, AutoTransform, T5-Review, CodeBERT, GraphCodeBERT, CodeT5, and CodeAgent. As detailed in Table 6, these approaches exhibit varied performance across different datasets. Notably, CodeAgent demonstrates a remarkable performance, particularly in the T5-Review dataset, where it achieves the highest EP of 37.6%. This is a significant improvement over other methods, underscoring the effectiveness of CodeAgent in handling complex code revision tasks. Additionally, with an average EP of 31.6%, CodeAgent consistently outperforms its counterparts, positioning itself as a leading solution in automated code revision. The contribution of CodeAgent to the field of automated code revision is noteworthy. Its ability to excel in the T5-Review dataset, a challenging benchmark, indicates a sophisticated understanding and handling of nuanced bugs. Moreover, its overall average performance surpasses that of other state-of-the-art models, highlighting its robustness and reliability. These results suggest that CodeAgent outperforms the state-of-the-art approaches in fixing buggy codes on metric on average.
| Approach | Average | |||
| EP | EP | EP | EP | |
| Trans-Review | -1.1% | -16.6% | -151.2% | -56.3% |
| AutoTransform | 49.7% | 29.9% | 9.7% | 29.8% |
| T5-Review | -14.9% | -71.5% | 13.8% | -24.2% |
| CodeBERT | 49.8% | -75.3% | 22.3% | -1.1% |
| GraphCodeBERT | 50.6% | -80.9% | 22.6% | -2.6% |
| CodeT5 | 41.8% | -67.8% | 25.6% | -0.1% |
| CodeAgent | 42.7% | 14.4% | 37.6% | 31.6% |
6 Related Work
Automating Code Review Activities Our work contributes to automating code review activities, focusing on detecting source code vulnerabilities and maintaining code consistency. Related studies include Hellendoorn et al. [15], who addressed code change anticipation, and Siow et al. [29], who introduced CORE for code modification semantics. Hong et al. [18] proposed COMMENTFINDER for comment suggestions, while Tufano et al. [35] and Li et al. [23] developed tools for code review automation using models like T5CR and CodeReviewer, respectively. Recently, Lu et al. [25] incorporated large language models for code review, enhancing fine-tuning techniques.
Collaborative AI Collaborative AI, involving AI systems working towards shared goals, has seen advancements in multi-agent LLMs [31; 28], focusing on collective thinking, conversation dataset curation [39; 21], and sociological phenomenon exploration [27]. Research by Akata et al. [1] and Cai et al. [5] further explores LLM cooperation and efficiency. However, there remains a gap in integrating these advancements with structured software engineering practices [21; 28], a challenge our approach addresses by incorporating advanced human processes in multi-agent systems. For a complete overview of related work, please refer to our Appendix-Section A.
7 Conclusion
This paper introduces CodeAgent, a novel multi-agent-based framework designed to automate code reviews, leveraging an innovative QA-Checker system to maintain focus and alignment with the review’s objectives. CodeAgent demonstrates promising capabilities in detecting vulnerabilities, ensuring consistency between code changes and commit messages, and improving uniformity of code style. Our CodeAgent outperforms existing state-of-the-art solutions. By considering the specific characteristics of the code review process and incorporating the human-like conversational framework, CodeAgent significantly enhances efficiency and accuracy. Finally, we believe that our paper opens a new avenue for future software development collaboration practice and research.
8 Impact Statement
The adoption of our multi-agent-based code review framework promises to revolutionize how software development teams ensure code quality. By automating and enhancing the code review process, developers can allocate more time to creative and value-added tasks, leading to increased productivity and a higher standard of software craftsmanship. We expect that the adaptive learning capabilities of our approach contribute to a sustainable and evolving solution, capable of addressing the dynamic challenges in modern software development and machine learning.
Appendix A Complete Related Work
Automating Code Review Activities Our focus included detecting source code vulnerabilities, ensuring style alignment, and maintaining commit message and code consistency. Other studies explore various aspects of code review. Hellendoorn et al. [15] addressed the challenge of anticipating code change positions. Siow et al. [29] introduced CORE, employing multi-level embeddings for code modification semantics and retrieval-based review suggestions. Hong et al. [18] proposed COMMENTFINDER, a retrieval-based method for suggesting comments during code reviews. Tufano et al. [35] designed T5CR with SentencePiece, enabling work with raw source code without abstraction. Li et al. [23] developed CodeReviewer, focusing on code diff quality, review comment generation, and code refinement using the T5 model. Recently, large language models have been incorporated; Lu et al. [25] fine-tuned LLama with prefix tuning for LLaMA-Reviewer, using parameter-efficient fine-tuning and instruction tuning in a code-centric domain.
Collaborative AI Collaborative AI refers to artificial intelligent systems designed to achieve shared goals with humans or other AI systems. Previous research extensively explores the use of multiple LLMs in collaborative settings, as demonstrated by Talebirad et al. [31] and Qian et al. [28]. These approaches rely on the idea that inter-agent interactions enable LLMs to collectively enhance their capabilities, leading to improved overall performance. The research covers various aspects of multi-agent scenarios, including collective thinking, conversation dataset curation, sociological phenomenon exploration, and collaboration for efficiency. Collective thinking aims to boost problem-solving abilities by orchestrating discussions among multiple agents. Researchers like Wei et al. [39] and Li et al. [21] have created conversational datasets through role-playing methodologies. Sociological phenomenon investigations, such as Park et al. [27]’s work, involve creating virtual communities with rudimentary language interactions and limited cooperative endeavors. In contrast, Akata et al. [1] scrutinized LLM cooperation through orchestrated repeated games. Collaboration for efficiency, proposed by Cai et al. [5], introduces a model for cost reduction through large models as tool-makers and small models as tool-users. Zhang et al. [42] established a framework for verbal communication and collaboration, enhancing overall efficiency. However, Li et al. [21] and Qian et al. [28], presenting a multi-agent framework for software development, primarily relied on natural language conversations, not standardized software engineering documentation, and lacked advanced human process management expertise. Challenges in multi-agent cooperation include maintaining coherence, avoiding unproductive loops, and fostering beneficial interactions. Our approach emphasizes integrating advanced human processes, like code review in software maintenance, within multi-agent systems.
Appendix B Experimental Details
In our work, the maximum number of conversation rounds is set as 10.
B.1 Role Definition
Six roles are defined as shown in Figure 5.
Apart from that, for the QA-checker in CodeAgent, we define an initial prompt for it, which is shown as follows:
B.2 Execute Time Across Languages
As depicted in the data, we observe a significant trend in the average execution time for code reviews in CodeAgent across various programming languages. The analysis includes nine languages: Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. For each language, the average execution time of code reviews for both merged and closed pull requests (PRs) is measured. The results, presented in Figure 6, indicate that, on average, the execution time for merged PRs is longer than that for closed PRs by approximately 44.92 seconds. This considerable time difference can be attributed to several potential reasons. One primary explanation is that merged PRs likely undergo a more rigorous and detailed review process. They are intended to be integrated into the main codebase, and as such, contributors might be requested to update their commits in the PRs more frequently to adhere to the project’s high-quality standards. On the other hand, closed PRs, which are not meant for merging, might not require such extensive review processes, leading to shorter review times on average, which may also be the reason they are not merged into main projects.
Appendix C Comparative Analysis of QA-Checker AI System and Recursive Self-Improvement Systems
In this section, we will delve into the differences between QA-Checker and self-improvement systems [17], and underscore the importance of the QA-Checker in role conversations.
C.1 Comparison Table
We begin with a comparative overview presented in Table 7.
| Feature/System | QA-Checker AI System | Recursive Self-Improvement System |
| Application Focus |
Specialized for QA tasks with precise task execution |
Broad scope, covering various dimensions like software development and learning algorithms |
| Learning Mechanism |
Advanced optimization techniques for iterative improvement in QA |
Multi-level learning: learning, meta-learning, and recursive self-improvement |
| Scope of Improvement |
Focused on individual capability in specific QA tasks |
Enhances the entire system, including multi-agent interactions and communication protocols |
| Experience Integration |
Based on mathematical models to optimize answer quality |
Utilizes experiences from past projects to improve overall performance |
C.2 Differences and Implications
The key differences between these systems lie in their application scope, learning mechanisms, and improvement scopes. The QA-Checker is highly specialized, focusing on QA tasks with efficiency and precision. In contrast, recursive self-improvement systems boast a broader application range and adaptability, integrating experiences from diverse projects for systemic improvements.
C.3 Importance of QA-Checker in Role Conversations
In the context of role conversations, the QA-Checker plays a pivotal role. Its specialized nature makes it exceptionally adept at handling specific conversational aspects, such as accuracy, relevance, and clarity in responses. This specialization is crucial in domains where the quality of information is paramount, ensuring that responses are not only correct but also contextually appropriate and informative.
Furthermore, the efficiency of the QA-Checker in refining responses based on advanced optimization techniques makes it an invaluable tool in dynamic conversational environments. It can quickly adapt to the nuances of a conversation, providing high-quality responses that are aligned with the evolving nature of dialogue.
C.4 Conclusion
While recursive self-improvement systems offer broad adaptability and systemic learning, the QA-Checker stands out in its specialized role in QA tasks, particularly in role conversations. Its focused approach to improving answer quality and its efficiency in handling conversational nuances make it an essential component in AI-driven communication systems.
Appendix D Capabilities Analysis between CodeAgent and Other Methods
Compared to open-source baseline methods such as AutoGPT and autonomous agents such as ChatDev and MetaGPT, CodeAgent offers functions for code review tasks: consistency analysis, vulnerability analysis, and format analysis. As shown in Table 8, our CodeAgent encompasses a wide range of abilities to handle complex code review tasks efficiently. Incorporating the QA-Checker self-improved module can significantly improve the conversation generation between agents and contribute to the improvement of code review. Compared to COT, the difference and the advantages of CodeAgent with QA-Checker are shown in Section C.
Appendix E Dataset
Previous Dataset
As shown in [44], our study incorporates three distinct datasets for evaluating the performance of CodeAgent: , , and . , compiled by Tufano et al. [35], derives from Gerrit and GitHub projects, excluding noisy or overly lengthy comments and review data with new tokens in revised code not present in the initial submission. , collected by Thongtanunam et al. [32] from three Gerrit repositories, comprises only submitted and revised codes without review comments. Lastly, , gathered by Tufano et al. [34] from Java projects on GitHub, filters out noisy, non-English, and duplicate comments. These datasets are employed for Code Revision Before Review (CRB) and Code Revision After Review (CRA) tasks, with the exception of for CRA and Review Comment Generation (RCG) due to its lack of review comments.
New Dataset Design and Collection
To enhance our model evaluation and avoid data leakage, we curated a new dataset, exclusively collecting data from repositories created after April 2023. This approach ensures the evaluation of our CodeAgent model on contemporary and relevant data, free from historical biases. The new dataset is extensive, covering a broad spectrum of software projects across nine programming languages.
Dataset Description
Our dataset, illustrated in Fig. 8, encapsulates a detailed analysis of consistency and format detection in software development, spanning various programming languages. It includes CA (consistency between commit and commit message (See Table 1)) and FA (format consistency between commit and original (See Table 1)) data, segmented into positive and negative samples based on the merged and closed status of pull requests. For example, in Python, the dataset comprises 254 merged and 35 closed negative CA samples, alongside 803 merged and 213 closed positive CA samples, with corresponding distributions for other languages like Java, Go, C++, and more. Similarly, the FA data follows this pattern of positive and negative samples across languages. Figure 7 graphically represents this data, highlighting the distribution and comparison of merged versus closed samples in both CA and FA categories for each language. This comprehensive dataset, covering over 3,545 commits and nearly 2,933 pull requests from more than 180 projects, was meticulously compiled using a custom crawler designed for GitHub API interactions, targeting post-April 2023 repositories to ensure up-to-date and diverse data for an in-depth analysis of current software development trends.
| Dataset Statistics | #Train | #Valid | #Test |
| Trans-Review | 13,756 | 1,719 | 1,719 |
| AutoTransform | 118,039 | 14,750 | 14,750 |
| T5-Review | 134,239 | 16,780 | 16,780 |
Appendix F Key Factors Leading to Vulnerabilities
The following table outlines various key factors that can lead to vulnerabilities in software systems, along with their descriptions. These factors should be carefully considered and addressed to enhance the security of the system.
| No. | Vulnerability Factor | Description |
| 1 |
Insufficient Input Validation |
Check for vulnerabilities like SQL injection, Cross-Site Scripting (XSS), and command injection in new or modified code, especially where user input is processed. |
| 2 |
Buffer Overflows |
Particularly in lower-level languages, ensure that memory management is handled securely to prevent overflows. |
| 3 |
Authentication and Authorization Flaws |
Evaluate any changes in authentication and authorization logic for potential weaknesses that could allow unauthorized access or privilege escalation. |
| 4 |
Sensitive Data Exposure |
Assess handling and storage of sensitive information like passwords, private keys, or personal data to prevent exposure. |
| 5 |
Improper Error and Exception Handling |
Ensure that errors and exceptions are handled appropriately without revealing sensitive information or causing service disruption. |
| 6 |
Vulnerabilities in Dependency Libraries or Components |
Review updates or changes in third-party libraries or components for known vulnerabilities. |
| 7 |
Cross-Site Request Forgery (CSRF) |
Verify that adequate protection mechanisms are in place against CSRF attacks. |
| 8 |
Unsafe Use of APIs |
Check for the use of insecure encryption algorithms or other risky API practices. |
| 9 |
Code Injection |
Look for vulnerabilities related to dynamic code execution. |
| 10 |
Configuration Errors |
Ensure that no insecure configurations or settings like open debug ports or default passwords have been introduced. |
| 11 |
Race Conditions |
Analyze for potential data corruption or security issues arising from race conditions. |
| 12 |
Memory Leaks |
Identify any changes that could potentially lead to memory leaks and resource exhaustion. |
| 13 |
Improper Resource Management |
Check resource management, such as proper closure of file handles or database connections. |
| 14 |
Inadequate Security Configurations |
Assess for any insecure default settings or unencrypted communications. |
| 15 |
Path Traversal and File Inclusion Vulnerabilities |
Examine for risks that could allow unauthorized file access or execution. |
| 16 |
Unsafe Deserialization |
Look for issues that could allow the execution of malicious code or tampering with application logic. |
| 17 |
XML External Entity (XXE) Attacks |
Check if XML processing is secure against XXE attacks. |
| 18 |
Inconsistent Error Handling |
Review error messages to ensure they do not leak sensitive system details. |
| 19 |
Server-Side Request Forgery (SSRF) |
Analyze for vulnerabilities that could be exploited to attack internal systems. |
| 20 |
Unsafe Redirects and Forwards |
Check for vulnerabilities leading to phishing or redirection attacks. |
| 21 |
Use of Deprecated or Unsafe Functions and Commands |
Identify usage of any such functions and commands in the code. |
| 22 |
Code Leakages and Hardcoded Sensitive Information |
Look for hardcoded passwords, keys, or other sensitive data in the code. |
| 23 |
Unencrypted Communications |
Verify that data transmissions are securely encrypted to prevent interception and tampering. |
| 24 |
Mobile Code Security Issues |
For mobile applications, ensure proper handling of permission requests and secure data storage. |
| 25 |
Cloud Service Configuration Errors |
Review any cloud-based configurations for potential data leaks or unauthorized access. |
Appendix G Data Leakage Statement
As the new dataset introduced in Section E, the time of the collected dataset is after April 2023, avoiding data leakage while we evaluate CodeAgent on codeData dataset.
Appendix H Algorithmic Description of CodeAgent Pipeline with QA-Checker
This algorithm demonstrates the integration of QA-Checker within the CodeAgent pipeline, employing mathematical equations to describe the QA-Checker’s iterative refinement process.
In this algorithm, represents the quality assessment function of the QA-Checker, which evaluates the relevance and accuracy of the answer to the question . If the quality score is below a predefined threshold, the QA-Checker intervenes by generating an additional instruction to refine the question, prompting a more accurate response in the next iteration.
Appendix I Detailed Performance of CodeAgent in Various Languages on VA task
In our comprehensive analysis using CodeAgent, as detailed in Table 11, we observe a diverse landscape of confirmed vulnerabilities across different programming languages. The table categorizes these vulnerabilities into ‘merged’ and ‘closed’ statuses for languages such as Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. A significant finding is a markedly high number of ‘merged’ vulnerabilities in Python, potentially reflective of its extensive application or intrinsic complexities leading to security gaps. Conversely, languages like Go, Ruby, and C exhibit notably lower counts in both categories, perhaps indicating lesser engagement in complex applications or more robust security protocols. Table 11 that the ‘closed’ category consistently presents lower vulnerabilities than ‘merged’ across most languages, signifying effective resolution mechanisms. However, an exception is noted in C, where ‘closed’ counts surpass those of ‘merged’, possibly indicating either delayed vulnerability identification or efficient mitigation strategies. Remarkably, the Rate is generally observed to be higher than Rate across the languages, exemplifying a significant reduction in vulnerabilities post-resolution. For example, Python demonstrates a Rate of 14.00% against a higher Rate of 18.16%. This trend is consistent in most languages, emphasizing the importance of proactive vulnerability management. The Rate, representing the proportion of confirmed vulnerabilities against the total of both merged and closed items, further elucidates this point, with C++ showing the highest Rate at 16.49%. These insights not only underline the diverse vulnerability landscape across programming languages but also highlight the adeptness of CodeAgent in pinpointing and verifying vulnerabilities in these varied contexts.
| CodeAgent | Python | Java | Go | C++ | JavaScript | C | C# | PHP | Ruby |
| merged (total#) | 1,057 | 287 | 133 | 138 | 280 | 114 | 206 | 173 | 202 |
| merged (confirmed#) | 148 | 17 | 11 | 19 | 34 | 9 | 21 | 28 | 20 |
| Rate | 14.00% | 5.92% | 8.27% | 13.77% | 12.14% | 7.89% | 10.19% | 16.18% | 9.90% |
| closed (total#) | 248 | 97 | 74 | 56 | 112 | 146 | 62 | 105 | 55 |
| closed (confirmed#) | 45 | 10 | 5 | 13 | 16 | 26 | 7 | 15 | 5 |
| Rate | 18.16% | 10.31% | 6.76% | 23.2% | 14.29% | 17.81% | 11.29% | 14.29% | 9.09% |
| Total number (#) | 1,305 | 384 | 207 | 194 | 392 | 260 | 268 | 278 | 257 |
| Total confirmed (#) | 193 | 27 | 16 | 32 | 50 | 35 | 28 | 43 | 25 |
| Rate | 14.79% | 7.03% | 7.73% | 16.49% | 12.76% | 13.46% | 10.45% | 14.47% | 9.73% |
Appendix J More detailed experimental results on CA and FA tasks
Detailed experimental results of CA are shown in Figure 9 and Figure 10. Detailed experimental results of FA are shown in Figure 11 and Figure 12.
Appendix K Case Study
K.1 Performance on 9 languages
K.2 Difference of CodeAgent-3.5 and CodeAgent-4.0
CodeAgent-3.5 and CodeAgent-4.0 in this paper has no difference in general code review, however, as CodeAgent-4.0 is more powerful in processing long input sequences and logic reasoning. As shown in Figure 22, we take one example of consistency detection between commit and commit message and find that CodeAgent-4.0 diffs from CodeAgent-3.5 in the detailed explanation. CodeAgent-3.5 output a report with 15k lines while CodeAgent-4.0 outputs a report with more than 17.7k lines. Detailed data is shown in https://zenodo.org/records/10607925.
Appendix L Ablation study
In this section, we evaluate the performance of different parts in CodeAgent in vulnerability analysis. CodeAgent is based on chain-of-thought (COT) and large language model (a.k.a. ChatGPT). As shown in Section 5.1, CodeAgent outperforms baselines (a.k.a. CodeBERT, ChatGPT-3.5, ChatGPT-4.0) across 9 different languages. The performance mainly comes from the combination of COT and QA-Checker. Thus, we design an additional version called CodeAgent , which means CodeAgent without QA-Checker. Then, we use CodeAgent to do vulnerability analysis and compare with CodeAgent. We first discuss about the result of CodeAgent and then discuss about comparison between CodeAgent and CodeAgent .
Overview of Vulnerabilities in CodeAgent
Table 12 presents the findings of CodeAgent , a variant of the original CodeAgent, in identifying vulnerabilities across different programming languages. The table showcases the number of ‘merged’ and ‘closed’ vulnerabilities in languages such as Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. Notably, Python leads in the ‘merged’ category with a total of 1,057 cases, of which 140 are confirmed, yielding a Rate of 13.25%. In contrast, languages like Go and Ruby show lower vulnerability counts in both ‘merged’ and ‘closed’ categories. The table also includes Rate and Rate, providing insights into the effectiveness of vulnerability management across these languages.
Detailed Comparison between CodeAgent and CodeAgent
Comparing the findings in Table 12 with those in Table 11, we observe some notable differences in vulnerability detection by CodeAgent and CodeAgent . While the overall trend of higher ‘merged’ vulnerabilities in Python and lower counts in Go and Ruby remains consistent, Table 12 shows a slight reduction in the Rate for most languages, suggesting a more conservative confirmation approach in CodeAgent . Similarly, Rate and Rate values in Table 12 generally indicate a lower proportion of confirmed vulnerabilities compared to Table 11, reflecting potentially different criteria or efficacy in vulnerability assessment. These variations highlight the impact of QA-Checker in CodeAgent.
| CodeAgent | Python | Java | Go | C++ | JavaScript | C | C# | PHP | Ruby |
| merged (total#) | 1,057 | 287 | 133 | 138 | 280 | 114 | 206 | 173 | 202 |
| merged (confirmed#) | 140 | 17 | 10 | 12 | 28 | 9 | 21 | 28 | 17 |
| Rate | 13.25% | 5.92% | 7.52% | 8.70% | 10.00% | 7.89% | 10.19% | 16.18% | 8.42% |
| closed (total#) | 248 | 97 | 74 | 56 | 112 | 146 | 62 | 105 | 55 |
| closed (confirmed#) | 36 | 9 | 5 | 12 | 16 | 26 | 7 | 15 | 5 |
| Rate | 14.52% | 9.28% | 6.76% | 21.43% | 14.29% | 17.81% | 11.29% | 14.29% | 9.09% |
| Total number (#) | 1,305 | 384 | 207 | 194 | 392 | 260 | 268 | 278 | 257 |
| Total confirmed (#) | 176 | 26 | 15 | 24 | 44 | 35 | 28 | 43 | 22 |
| Rate | 13.49% | 6.77% | 7.25% | 12.37% | 11.22% | 13.46% | 10.45% | 15.47% | 8.56% |
Appendix M Tool
We develop a website for CodeAgent, which is shown in Figure 23, and it is also accessable by visiting following link: