### v4

# Haoye Tian

Abstract
Code review, which aims at ensuring the overall quality and reliability of software, is a cornerstone of software development. Unfortunately, while crucial, Code review is a labor-intensive process that the research community is looking to automate. Existing automated methods rely on single input-output generative models and thus generally struggle to emulate the collaborative nature of code review. This work introduces CodeAgent, a novel multi-agent Large Language Model (LLM) system for code review automation. CodeAgent incorporates a supervisory agent, QA-Checker, to ensure that all the agents’ contributions address the initial review question. We evaluated CodeAgent on critical code review tasks: (1) detect inconsistencies between code changes and commit messages, (2) identify vulnerability introductions, (3) validate code style adherence, and (4) suggest code revisions. The results demonstrate CodeAgent’s effectiveness, contributing to a new state-of-the-art in code review automation. Our data and code are publicly available (https://github.com/Code4Agent/codeagent).
[…]
CodeAgent: Autonomous Communicative Agents for Code Review
1 Introduction
Code review Bacchelli and Bird (2013); Bosu and Carver (2013); Davila and Nunes (2021) implements a process wherein software maintainers examine and assess code contributions to ensure quality and adherence to coding standards, and identify potential bugs or improvements. In recent literature, various approaches Tufano et al. (2021, 2022) have been proposed to enhance the performance of code review automation. Unfortunately, major approaches in the field ignore a fundamental aspect: the code review process is inherently interactive and collaborative Bacchelli and Bird (2013). Instead, they primarily focus on rewriting and adapting the submitted code Watson et al. (2022); Thongtanunam et al. (2022); Staron et al. (2020). In this respect, an effective approach should not only address how to review the submitted code for some specific needs (e.g., vulnerability detection). Still, other non-negligible aspects of code review should also be considered, like detecting issues in code formatting or inconsistencies in code revisions. However, processing multiple sub-tasks requires interactions among different staff members in a real code review scenario, which makes it chellenging to design a bot that performs code review automatically.
Agent-based systems are a newly emerging paradigm and a computational framework in which autonomous entities (aka agents) interact with each other Li et al. (2023a); Qian et al. (2023); Hong et al. (2023) to perform a task. Agent-based approaches have been proposed to address a spectrum of software engineering tasks Qian et al. (2023); Zhang et al. (2024), moving beyond the conventional single input-output paradigm due to their remarkable ability to simulate and model complex interactions and behaviors within dynamic environments Xi et al. (2023); Wang et al. (2023). Recently, with multi-agent systems, the strengths of various agents are utilized to simulate human-like decision-making processes Du et al. (2023); Liang et al. (2023); Park et al. (2023), improving performance in many tasks Chen et al. (2023); Li et al. (2023b); Hong et al. (2023). This paradigm fits well with the challenge of code review, which involves multiple reviewers interacting based on their diverse skills and roles to achieve a thorough review of code.
This paper. Drawing from the success of agent-based collaboration, we propose a multi-agent-based framework CodeAgent to simulate the dynamics of a collaborative team engaged in the code review process, incorporating various roles such as code change authors, reviewers, and decision makers. In particular, as a major contribution towards CodeAgent, we address the challenge of prompt drifting Zheng et al. (2024); Yang et al. (2024), commonly found in multi-agent systems and Chain-of-Thought (CoT) reasoning. This issue, characterized by conversations that stray from the main topic, highlights the need for strategies to keep discussions on track and coherent Greyling (2023); Chae et al. (2023). This drift, often triggered by the model-inspired tangents or the randomness of Large Language Models (LLMs), necessitates the integration of a supervisory agent. We employ an agent named QA-Checker (for "Question-Answer Checker") that monitors the flow of conversation, ensuring that questions and responses remain relevant and on track, thus maintaining the dialogue’s intended objective. Such an agent not only refines queries but also realigns answers to match the original intent, employing a systematic approach grounded in a mathematical framework.
To evaluate the performance of CodeAgent, we first assess its effectiveness for typical review objectives such as finding vulnerabilities 4.1 and validating the consistency and alignment of the code format 4.2. We also compare CodeAgent with state-of-the-art generic and code-specific language models Feng et al. (2020); OPENAI (2022). Finally, we assess the performance of CodeAgent compared to the state of the art for code revision suggestions Tufano et al. (2021); Thongtanunam et al. (2022); Tufano et al. (2022). Since each of these related works presents a specific dataset, we also employ them toward a fair comparison. At the same time, we also collect pull requests from GitHub, featuring an extensive array of commits, messages, and comments to evaluate advanced capabilities.The experimental results display that CodeAgent significantly outperforms the state-of-the-art, achieving a 41% increase in hit rate for detecting vulnerabilities. CodeAgent also excels in consistency checking and format alignment, outperforming the target models. Finally, CodeAgent showcases its robustness for code revision by demonstrating superior average edit progress.
We summarize our contributions as follows:
  • To the best of our knowledge, we are the first to propose an autonomous agent-based system for practical code review in the field of software maintenance.
  • We build a new dataset with about than 3 500 real-world code changes and commit messages. This dataset, which includes all relevant files and details in a standalone manner, is useful for assessing advanced code review tasks such as vulnerability detection, code style detection, and code revision suggestions.
  • We demonstrate the effectiveness of the QA-Checker. This agent monitors the conversation flow to ensure alignment with the original intent, addressing the common prompt drifting issues in multi-agent systems.
Experimental evaluation highlights the performance of CodeAgent: In terms of vulnerability detection, CodeAgent outperforms GPT-4 and CodeBERT by 3 to 7 percentage points in terms of the number of vulnerabilities detected. In terms of format alignment, CodeAgent outperforms ReAct by about 14% in terms of recall of inconsistency detection. On the code revision task, CodeAgent surpasses the state of the art in the software engineering literature: in terms of the Edit Progress metric Zhou et al. (2023), CodeAgent offers an average performance improvement of about 30%.
2 CodeAgent
This section details the methodology behind our novel CodeAgent framework. We discuss tasks and definition in Sec 2.1, pipeline in Section 2.2, defined role cards in Section 2.3, and the design of the QA-Checker in Sec 2.4.
Figure 1: A Schematic diagram of role data cards of simulated code review team and their conversations within CodeAgent. We have six characters in CodeAgent across four phases, including “Basic Info Sync", “Code Review", “Code Alignment", and “Document". Code review is a kind of collaboration work, where we design conversations between every two roles for every step to complete the task.
Figure 2: CodeAgent’s pipeline/scenario of a full conversation during the code review process among different roles. “Basic Info Sync” demonstrates the basic information confirmation by the CEO, CTO, and Coder; “Code Review” shows the actual code review process; “Code Alignment” illustrates the potential code revision; and “Document” represents the summarizing and writing conclusion for all the stakeholders. All the conversations are being ensured by the Quality Assurance checker until they reach the maximum dialogue turns or meet all the requirements.
2.1 Tasks
We define CA, VA, FA, and CR in as following:
CA Zhang et al. (2022): Consistency analysis between code change and commit message; the task is to detect cases where the commit message accurate describes (in natural language) the intent of code changes (in programming language).
VA Braz et al. (2022): Vulnerability analysis; the task is to identify cases where the code change introduces a vulnerability in the code.
FA Han et al. (2020): Format consistency analysis between commit and original files; the task is to validate that the code change formatting style is not aligned with the target code.
CR Zhou et al. (2023): Code revision; this task attempts to automatically suggest rewrites of the code change to address any issue discovered.
2.2 Pipeline
We defined six characters and four phases for the framework. The roles of the characters are illustrated in Figure 1. Each phase contains multiple conversations, and each conversation happens between agents. The four phases consist of 1) Basic Info Sync, containing the roles of chief executive officer (CEO), chief technology officer (CTO), and Coder to conduct modality and language analysis; 2) Code Review, leveraging the Coder and Reviewer for actual code review (i.e., target sub-tasks); 3) Code Alignment, supporting the Coder and Reviewer to correct the commit by code revision or suggestion feedback to the author; and 4) Document, finalizing by synthesizing the opinions of the CEO, CPO, Coder, and Reviewer to give out the final comments. In addition to six defined roles, the proposed architecture of CodeAgent consists of phase-level and conversation-level components. The waterfall model breaks the code review process at the phase level into four sequential phases. At the conversation level, each phase is divided into atomic conversations. These atomic conversations involve task-oriented role-playing between two agents, promoting collaborative communication. One agent works as an instructor and the other as an assistant. Communication follows an instruction-following style, where agents interact to accomplish a specific subtask within each conversation, and each conversation is supervised by QA-Checker. QA-Checker is used to align the consistency of questions and answers between the instructor and the assistant in a conversation to avoid digression. QA-Checker will be introduced in Section 2.4.
Figure 2 shows an illustrative example of the CodeAgent pipeline. CodeAgent receives the request to do the code review with the submitted commit, commit message, and original files. In the first phase, CEO, CPO, and Coder will cooperate to recognize the modality of input (for example, document, code) and language (for example, Python, Java and Go). In the second phase, with the help of Coder, Reviewer will write an analysis report on consistency analysis, vulnerability analysis, format analysis and suggestions for code revision. Then, in the third phase, according to analysis reports, Coder will align or revise the code if it finds incorrect snippets with the help of Reviewer. $Coder$ cooperates with CTO and CEO to summarize the document and codes about the whole code review in the final phase.
2.3 Role Card Definition
As shown in Figure 1, we define six characters in our simulation system (CodeAgent), including User, CEO, CPO, CTO, Reviewer, Coder, and they are defined for different specific tasks.
All tasks are processed by the collaborative work of two agents in their multi-round conversations. For example, as a role Reviewer, her responsibility is to do the code review for given codes and files in three aspects (tasks CA, VA, and FA in Sec 2.1) and provide a detailed description of observation. Reviewer’s code review activity is under the assistance with Coder as shown in Figure 2. Meanwhile, with the Reviewer’s assistance, Coder can process the code revision as shown in the ‘Revised codes’ part in the Coder card in Figure 1. Apart from Reviewer, Coder also cooperates with CTO and CEO in the simulated team.
Each role and conversation, input and output of each conversation is designed in Figure 1. Further information about role definition details is provided in our Appendix-Section C.1.
2.4 Self-Improving CoT with QA Checker
Figure 3: This diagram shows the architecture of our designed Chain-of-Thought (CoT): Question-Answer Checker (QA-Checker).
QA-Checker is an instruct-driven agent, designed to fine-tune the question inside a conversation to drive the generated answer related to the question. As shown in Figure 3, the initial question (task instruction) is represented as $q_{0}$, and the first answer of the conversation between $Reviewer$ and $Coder$ is represented as $a_{0}$. If QA-Checker identifies that $a_{0}$ is inappropriate for $q_{0}$, it generates additional instructions attached to the original question (task instruction) and combines them to ask agents to further generate a different answer. The combination in Figure 3 is defined as $q_{1}=CB(q_{0}+aai_{0})$, where $aai_{0}$ is the additional instruction attached. The conversation between two agents is held until the generated answer is judged as appropriate by QA-Checker or it reaches the maximum number of dialogue turns.
Theoretical Analysis of QA-Checker in Dialogue Refinement
The QA-Checker is an instruction-driven agent, crucial in refining questions and answers within a conversation to ensure relevance and precision. Its operation can be understood through the following lemma and proof in Appendix A.
3 Experimental Setup
We evaluate the performance of CodeAgent in various qualitative and quantitative experiments in nine programming languages, on four metrics. In this section, we will discuss experimental settings, including datasets, metrics, and baselines. For more information, please see Appendix C.
3.1 Datasets
To conduct a fair and reliable comparison for the code revision task, we employ the same datasets (i.e., Trans-Reviewdata, AutoTransformdata, and T5-Reviewdata) as the state-of-the-art study Zhou et al. (2023). Furthermore, we collect and curate an additional dataset targeting the advanced tasks. Table 1 shows our new dataset which includes over 3,545 commits and 2,933 pull requests from more than 180 projects, spanning nine programming languages: Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. It focuses on consistency and format detection, featuring both positive and negative samples segmented by the merged and closed status of pull requests across various languages. The detailed information about the dataset can be seen in Appendix-Section F.
Table 1: Comparison of Positive and Negative Samples in CA and FA (CA and FA are defined in Section 2.1).
Samples CA FA
Merged Closed Merged Closed
Positive (consistency) 2,089 820 2,238 861
Negative (inconsistency) 501 135 352 94
3.2 Metrics
  • F1-Score and Recall. We utilized the F1-Score and recall to evaluate our method’s effectiveness on tasks CA and FA. The F1-Score, a balance between precision and recall, is crucial for distinguishing between false positives and negatives. Recall measures the proportion of actual positives correctly identified Hossin and Sulaiman (2015).
  • Edit Progress (EP). EP evaluates the improvement in code transitioning from erroneous to correct by measuring the reduction in edit distance between the original code and the prediction on task CR. A higher EP indicates better efficiency in code generation Dibia et al. (2022); Elgohary et al. (2021); Zhou et al. (2023).
  • Hit Rate (Rate) We also use hit rate to evaluate the rate of confirmed vulnerable issues out of the found issues by approaches on task VA.
Table 2: The number of vulnerabilities found by CodeAgent and other approaches. As described in Appendix-Section F, we have 3,545 items to evaluate. Ratecr represents the confirmed number divided by the number of findings while Rateca is the confirmed number divided by the total evaluated number. CodeAgent w/o indicates the version without QA-Checker.
CodeBERT GPT-3.5 GPT-4.0 COT ReAct CodeAgent CodeAgent w/o
Find 1,063 864 671 752 693 483 564
Confirm 212 317 345 371 359 449 413
Ratecr 19.94% 36.69% 51.42% 49.34% 51.80% 92.96% 73.23%
Rateca 5.98% 8.94% 9.73% 10.46% 10.13% 12.67% 11.65%
The values in gray (nn.nn) denote the greatest values for the confirmed number of vulnerabilities and the rates.
3.3 State-of-the-Art Tools and Models
Our study evaluates various tools and models for code revision and modeling. Trans-Review Tufano et al. (2021) employs src2abs for code abstraction, effectively reducing vocabulary size. AutoTransform Thongtanunam et al. (2022) uses Byte-Pair Encoding for efficient vocabulary management in pre-review code revisions. T5-Review Tufano et al. (2022) leverages the T5 architecture, emphasizing improvement in code review through pre-training on code and text data. In handling both natural and programming languages, CodeBERT Feng et al. (2020) adopts a bimodal approach, while GraphCodeBERT Guo et al. (2021) incorporates code structure into its modeling. CodeT5 Wang et al. (2021), based on the T5 framework, is optimized for identifier type awareness, aiding in generation-based tasks. Additionally, we compare these tools with GPT OPENAI (2022) by OpenAI, notable for its human-like text generation capabilities in natural language processing. Finally, we involve COT Wei et al. (2022) and ReAct Yao et al. (2022), of which COT is a method where language models are guided to solve complex problems by generating and following a series of intermediate reasoning steps and ReAct synergistically enhances language models by interleaving reasoning and action generation, improving task performance and interpretability across various decision-making and language tasks.
4 Experimental Result Analysis
This section discusses the performance of CodeAgent in the four tasks considered for our experiments. In Appendix Section E, we provide further analyses: we discuss the difference in the execution time of CodeAgent in different languages and perform a capability analysis between CodeAgent and recent approaches.
4.1 Vulnerability Analysis
Compared to CA and FA, VA is a more complex code review subtask, covering more than 25 different aspects (please see the Appendix-Section G), including buffer overflows, sensitive data exposure, configuration errors, data leakage, etc. Vulnerability analysis being a costly, time-consuming, resource-intensive and sensitive activity, only a low proportion of commits are labeled. We therefore propose a proactive method for data annotion: we execute CodeAgent on the 3,545 samples (covering nine languages) and manual verify the identified cases to build a ground truth. Then, we applied CodeBERT Feng et al. (2020) and GPT on the dataset with the task of vulnerability binary prediction.
Comparison
As shown in Table 2, CodeAgent successfully identified 483 potential vulnerabilities within a data set of 3,545 samples, with an impressive 449 of these finally confirmed as high-risk vulnerabilities [^1] The verification process involved a rigorous manual examination, extending beyond 120 working hours. Each sample being validated by at least 2 people: a researcher and an engineer. CodeBERT, a key pre-trained model for code-related tasks, with its parameters frozen for this experiment, initially identified 1,063 items as vulnerable, yet only 212 passed the stringent verification criteria. Similar trends were observed with GPT-3.5 and GPT-4.0, which confirmed 317 and 345 vulnerabilities out of 864 and 671 identified items, respectively. These outcomes are further quantified by the confirmation rates (Ratecr) of 19.94% for CodeBERT, 36.69% for GPT-3.5, and 51.42% for GPT-4.0, while CodeAgent demonstrated a remarkable Ratecr of 92.96%. Additionally, the analysis of confirmed vulnerabilities against all analyzed items (Rateca) yielded 5.98%, 8.94%, 9.73%, and 12.67% for CodeBERT, GPT-3.5, GPT-4.0, and CodeAgent, respectively. Evidently, Table 2 not only highlights CodeAgent’s high precision in identifying vulnerable commits but also reveals the progressive improvement from GPT-3.5 to GPT-4.0, likely due to the latter’s capacity to handle longer input sequences, with token limits of 4,096 and 32,768, respectively. The integration of sophisticated algorithms like CoT and QA-Checker in CodeAgent has significantly enhanced its capabilities in vulnerability detection, surpassing the individual input-output efficiencies of GPT and CodeBERT. Appendix-Sections D and  M highlight further details regarding the importance of the QA-checker. Moreover, more experimental results in 9 languages are accessible in Appendix-Section J.
In addition, the analysis of vulnerabilities identified by various models reveals interesting overlaps among the models. CodeBERT confirmed 212 vulnerabilities, whereas GPT-3.5, GPT-4.0, and CodeAgent confirmed 317, 345, and 449 vulnerabilities, respectively. Notably, the intersection of vulnerabilities confirmed by CodeBERT and GPT-3.5 is 169, indicating a substantial overlap in their findings. Similarly, the intersection between CodeBERT and GPT-4.0 is 170, while a larger overlap of 212 vulnerabilities is observed between GPT-3.5 and GPT-4.0. The combined intersection among CodeBERT, GPT-3.5, and GPT-4.0 is 137, underscoring the commonalities in vulnerabilities detected across these models. Further, the intersections of vulnerabilities confirmed by CodeBERT, GPT-3.5, and GPT-4.0 with CodeAgent are 212, 317, and 334, respectively, highlighting the comprehensive coverage and detection capabilities of CodeAgent.
Figure 4: Overlap of vulnerability detection by CodeBERT, GPT-3.5, GPT-4.0, and CodeAgent.
Ablation Study.
As shown in Table 2, we conducted an ablation study to evaluate the effectiveness of the QA-Checker in CodeAgent. Specifically, we created a version of our tool without the QA-Checker, referred to as CodeAgent w/o. We then compared this version to the full version of CodeAgent that includes the QA-Checker. The results demonstrate that CodeAgent w/o is substantially less effective in identifying vulnerable issues, yielding lower hit rates (Ratecr and Rateca). This reduction in performance highlights the critical role of the QA-Checker in enhancing CodeAgent’s overall effectiveness. More detailed information about the ablation study can be found in Appendix-Section M.
4.2 Consistency and Format Detection
In this section, we will discuss the performance of CodeAgent and baselines on metrics like the F1-Score and recall score of task CA and FA. For CA and FA, the dataset we have is shown in Table 1 and more detailed data information is shown in Figure 7 in Appendix.
Code Change and Commit Message Consistency Detection.
As illustrated in Table 3, we assess the efficacy of CodeAgent in detecting the consistency between code changes and commit messages, contrasting its performance with other prevalent methods like CodeBERT, GPT-3.5, and GPT-4.0. This evaluation specifically focuses on merged and closed commits in nine languages. In particular, CodeAgent exhibits remarkable performance, outperforming other methods in both merged and closed scenarios. In terms of Recall, CodeAgent achieved an impressive 90.11% for merged commits and 87.15% for closed ones, marking a considerable average improvement of 5.62% over the other models. Similarly, the F1-Score of CodeAgent stands at 93.89% for merged and 92.40% for closed commits, surpassing its counterparts with an average improvement of 3.79%. More comparable details in different languages are shown in Appendix-Section. K.
Table 3: Comparison of CodeAgent with other methods on merged and closed commits across 9 languages on CA task. ‘Imp’ represents the improvement.
Merged CodeBERT GPT-3.5 GPT-4.0 COT ReAct CodeAgent Imp (pp)
Recall 63.64 80.08 84.27 80.73 82.04 90.11 5.84
F1 75.00 87.20 90.12 87.62 88.93 93.89 3.77
Closed CodeBERT GPT-3.5 GPT-4.0 COT ReAct CodeAgent Imp (pp)
Recall 64.80 79.05 81.75 81.77 83.42 87.15 5.21
F1 77.20 87.35 89.61 89.30 89.81 92.40 3.35
Average CodeBERT GPT-3.5 GPT-4.0 COT ReAct CodeAgent Imp (pp)
Recall 64.22 79.57 83.01 81.25 82.73 88.63 5.62
F1 76.01 87.28 89.61 88.46 89.37 93.16 3.79
Format Consistency Detection.
In our detailed evaluation of format consistency between commits and original files, CodeAgent’s performance was benchmarked against established models like CodeBERT and GPT variants across nine different languages. This comparative analysis, presented in Table 4, was centered around pivotal metrics such as Recall and F1-Score. CodeAgent demonstrated a significant edge over the state-of-the-art, particularly in the merged category, with an impressive Recall of 89.34% and an F1-Score of 94.01%. These figures represent an average improvement of 10.81% in Recall and 6.94% in F1-Score over other models. In the closed category, CodeAgent continued to outperform, achieving a Recall of 89.57% and an F1-Score of 94.13%, surpassing its counterparts with an improvement of 15.56% in Recall and 9.94% in F1-Score. The overall average performance of CodeAgent further accentuates its superiority, with a Recall of 89.46% and an F1-Score of 94.07%, marking an average improvement of 13.39% in Recall and 10.45% in F1-Score. These results underscore CodeAgent’s exceptional capability in accurately detecting format consistency between commits and their original files.
Table 4: Comparison of CodeAgent with other methods on merged and closed commits across the 9 languages on FA task. ‘Imp’ represents the improvement.
Merged CodeBERT GPT-3.5 GPT-4.0 COT ReAct CodeAgent Imp (pp)
Recall 60.59 60.72 78.53 70.39 71.21 89.34 10.81
F1 74.14 74.88 87.07 80.69 82.18 94.01 6.94
Closed CodeBERT GPT-3.5 GPT-4.0 COT ReAct CodeAgent Imp (pp)
Recall 69.95 73.61 68.46 73.39 74.01 89.57 15.56
F1 80.49 84.19 80.16 83.65 83.90 94.13 9.94
Average CodeBERT GPT-3.5 GPT-4.0 COT ReAct CodeAgent Imp (pp)
Recall 65.27 67.17 73.50 71.89 72.61 89.46 15.96
F1 77.32 79.54 83.62 82.17 83.04 94.07 10.45
4.3 Code Revision
We evaluate the effectiveness of CodeAgent in revision suggestion (i.e., bug fixing) based on Edit Progress (EP) metric. We consider Trans-Review, AutoTransform, T5-Review, CodeBERT, GraphCodeBERT, CodeT5 as comparable state of the art. As detailed in Table 5, these approaches exhibit a varied performance across different datasets. In particular, CodeAgent shows remarkable performance in the T5-Review dataset, achieving the highest EP of 37.6%. This is a significant improvement over other methods, which underlines the effectiveness of CodeAgent in handling complex code revision tasks. Furthermore, with an average EP of 31.6%, CodeAgent consistently outperforms its counterparts, positioning itself as a leading solution in automated code revision. Its ability to excel in the T5-Review, a challenging benchmark data, indicates a strong capability to address complex bugs. In addition, its overall average performance surpasses other state-of-the-art models, highlighting its robustness and reliability.
Table 5: Experimental Results for the Code Revision (CR task) of CodeAgent and the state-of-the-art works. Bold indicates the best performers.
[…]
EP EP EP EP
Trans-Review -1.1% -16.6% -151.2% -56.3%
AutoTransform 49.7% 29.9% 9.7% 29.8%
T5-Review -14.9% -71.5% 13.8% -24.2%
CodeBERT 49.8% -75.3% 22.3% -1.1%
GraphCodeBERT 50.6% -80.9% 22.6% -2.6%
CodeT5 41.8% -67.8% 25.6% -0.1%
CodeAgent 42.7% 14.4% 37.6% 31.6%
5 Related Work
Automating Code Review Activities. Our work contributes to automating code review activities, focusing on detecting source code vulnerabilities and maintaining code consistency. Related studies include Hellendoorn et al. Hellendoorn et al. (2021), who addressed code change anticipation, and Siow et al. Siow et al. (2020), who introduced CORE for code modification semantics. Hong et al. Hong et al. (2022) proposed COMMENTFINDER for comment suggestions, while Tufano et al. Tufano et al. (2021) and Li et al. Li et al. (2022) developed tools for code review automation using models like T5CR and CodeReviewer, respectively. Recently, Lu et al. Lu et al. (2023) incorporated large language models for code review, enhancing fine-tuning techniques.
Collaborative AI. Collaborative AI, involving AI systems working towards shared goals, has seen advancements in multi-agent LLMs Talebirad and Nadiri (2023); Qian et al. (2023), focusing on collective thinking, conversation dataset curation Wei et al. (2023); Li et al. (2023a), and sociological phenomenon exploration Park et al. (2023). Research by Akata et al. Akata et al. (2023) and Cai et al. Cai et al. (2023) further explores LLM cooperation and efficiency. However, there remains a gap in integrating these advancements with structured software engineering practices Li et al. (2023a); Qian et al. (2023), a challenge our approach addresses by incorporating advanced human processes in multi-agent systems. For a complete overview of related work, please refer to our Appendix-Section B.
6 Conclusion
In this paper, we introduced CodeAgent, a novel multi-agent framework that automates code reviews. CodeAgent leverages its novel QA-Checker system to maintain focus on the review’s objectives and ensure alignment. Our experiments demonstrate CodeAgent’s effectiveness in detecting vulnerabilities, enforcing code-message consistency, and promoting uniform code style. Furthermore, CodeAgent outperforms existing state-of-the-art solutions in code revision suggestions. By incorporating human-like conversational elements and considering the specific characteristics of code review, CodeAgent significantly improves both efficiency and accuracy. We believe this work opens exciting new avenues for research and collaboration practices in software development.
Limitations
Firstly, the generalizability of the system across different software development environments or industries may require further validation and testing. While the system has shown promising results in the provided datasets, its applicability to other contexts remains uncertain without additional empirical evidence. This limitation suggests that the findings may not be fully transferable to all settings within the software development domain. Secondly, the baseline test used in the study might be insufficient. The current testing approach may not fully capture the system’s performance, particularly in edge cases or more complex scenarios. This could result in an overestimation of the system’s capabilities and an underestimation of its limitations. Further, more comprehensive testing is needed to establish a more robust baseline and to ensure that the system performs reliably across a wider range of conditions.
Ethics Statements
This study was conducted in compliance with ethical guidelines and standards for research. The research did not involve human participants, and therefore, did not require informed consent or ethical review from an institutional review board. All data used in this study were publicly available, and no personal or sensitive information was accessed or processed. The development and evaluation of the CodeAgent system were performed with a focus on transparency, reproducibility, and the potential positive impact on the software development community.
Appendix A Details of QA-Checker Algorithm
Lemma A.1.
Let $\mathcal{Q}(Q_{i},A_{i})$ denote the quality assessment function of the QA-Checker for the question-answer pair $(Q_{i},A_{i})$ in a conversation at the $i$-th iteration. Assume $\mathcal{Q}$ is twice differentiable and its Hessian matrix $H(\mathcal{Q})$ is positive definite. If the QA-Checker modifies the question $Q_{i}$ to $Q_{i+1}$ by attaching an additional instruction $aai_{i}$, and this leads to a refined answer $A_{i+1}$, then the sequence $\{(Q_{i},A_{i})\}$ converges to an optimal question-answer pair $(Q^{*},A^{*})$, under specific regularity conditions.
Proof.
The QA-Checker refines the question and answers using the rule:
$Q_{i+1}$ $=Q_{i}+aai_{i},$
$A_{i+1}$ $=A_{i}-\alpha H(\mathcal{Q}(Q_{i},A_{i}))^{-1}\nabla\mathcal{Q}(Q_{i},A_{i}),$
where $\alpha$ is the learning rate. To analyze convergence, we consider the Taylor expansion of $\mathcal{Q}$ around $(Q_{i},A_{i})$:
$\mathcal{Q}(Q_{i+1},A_{i+1})$ $\approx\mathcal{Q}(Q_{i},A_{i})+\nabla\mathcal{Q}(Q_{i},A_{i})$
$\cdot(Q_{i+1}-Q_{i},A_{i+1}-A_{i})$
$\quad+\frac{1}{2}(Q_{i+1}-Q_{i},A_{i+1}-A_{i})^{T}$
$H(\mathcal{Q}(Q_{i},A_{i}))(Q_{i+1}-Q_{i},A_{i+1}-A_{i}).$
Substituting the update rule and rearranging, we get:
$\mathcal{Q}(Q_{i+1},A_{i+1})$ $\approx\mathcal{Q}(Q_{i},A_{i})$
$-\alpha\nabla\mathcal{Q}(Q_{i},A_{i})^{T}H(\mathcal{Q}(Q_{i},A_{i}))^{-1}$
$\nabla\mathcal{Q}(Q_{i},A_{i})$
$+\frac{\alpha^{2}}{2}\nabla\mathcal{Q}(Q_{i},A_{i})^{T}H(\mathcal{Q}(Q_{i},A_{i}))^{-1}$
$\nabla\mathcal{Q}(Q_{i},A_{i}).$
For sufficiently small $\alpha$, this model suggests an increase in $\mathcal{Q}$, implying convergence to an optimal question-answer pair $(Q^{*},A^{*})$ as $i\to\infty$. The convergence relies on the positive definiteness of $H(\mathcal{Q})$ and the appropriate choice of $\alpha$, ensuring each iteration moves towards an improved quality of the question-answer pair. ∎
In practical terms, this lemma and its proof underpin the QA-Checker’s ability to refine answers iteratively. The QA-Checker assesses the quality of each answer concerning the posed question, employing advanced optimization techniques that are modeled by the modified Newton-Raphson method to enhance answer quality. This framework ensures that, with each iteration, the system moves closer to the optimal answer, leveraging both first and second-order derivatives for efficient and effective learning.
Further Discussion
The QA-Checker computes $\mathcal{Q}(Q_{i},A_{i})$ at each iteration $i$ and compares it to a predefined quality threshold $\tau$. If $\mathcal{Q}(Q_{i},A_{i})<\tau$, the QA-Checker generates an additional instruction $aai_{i}$ to refine the question to $Q_{i+1}=Q_{i}+aai_{i}$, prompting the agents to generate an improved answer $A_{i+1}$.
First, we assume that the quality assessment function $\mathcal{Q}(Q_{i},A_{i})$ is twice differentiable with respect to the question $Q_{i}$. This assumption is reasonable given the smooth nature of the component functions (relevance, specificity, and coherence) and the use of continuous word embeddings. Next, we apply the second-order Taylor approximation to $\mathcal{Q}(Q_{i+1},A_{i+1})$ around the point $(Q_{i},A_{i})$:
$\mathcal{Q}(Q_{i+1},A_{i+1})$ $\approx\mathcal{Q}(Q_{i},A_{i})+\nabla\mathcal{Q}(Q_{i},A_{i})^{T}\Delta Q_{i}$
$+\frac{1}{2}\Delta Q_{i}^{T}H(\mathcal{Q}(Q_{i},A_{i}))\Delta Q_{i}+R_{2}(\Delta Q_{i})$
where $\Delta Q_{i}=Q_{i+1}-Q_{i}$, $H(\mathcal{Q}(Q_{i},A_{i}))$ is the Hessian matrix of $\mathcal{Q}$ evaluated at $(Q_{i},A_{i})$, and $R_{2}(\Delta Q_{i})$ is the remainder term.
Assuming that the remainder term $R_{2}(\Delta Q_{i})$ is negligible and that the Hessian matrix is positive definite, we can approximate the optimal step $\Delta Q_{i}^{*}$ as:
$\Delta Q_{i}^{*}\approx-H(\mathcal{Q}(Q_{i},A_{i}))^{-1}\nabla\mathcal{Q}(Q_{i},A_{i}).$
Substituting this approximation into the Taylor expansion and using the fact that $Q_{i+1}=Q_{i}+\alpha\Delta Q_{i}^{*}$ (where $\alpha$ is the learning rate), we obtain:
$\mathcal{Q}(Q_{i+1},A_{i+1})\approx$ $\mathcal{Q}(Q_{i},A_{i})-\alpha\nabla\mathcal{Q}(Q_{i},A_{i})^{T}$
$\cdot H(\mathcal{Q}(Q_{i},A_{i}))^{-1}\nabla\mathcal{Q}(Q_{i},A_{i})$
$+\frac{\alpha^{2}}{2}\nabla\mathcal{Q}(Q_{i},A_{i})^{T}H(\mathcal{Q}(Q_{i},A_{i}))^{-1}$
$\cdot\nabla\mathcal{Q}(Q_{i},A_{i}).$
The assumptions of twice differentiability, negligible remainder term, and positive definite Hessian matrix provide a more solid foundation for the approximation in Lemma 3.1. For sufficiently small $\alpha$, this approximation suggests an increase in $\mathcal{Q}$, implying convergence to an optimal question-answer pair $(Q^{*},A^{*})$ as $i\to\infty$. The convergence relies on the positive definiteness of $H(\mathcal{Q})$ and the appropriate choice of $\alpha$, ensuring each iteration moves towards an improved quality of the question-answer pair.
The quality assessment function $\mathcal{Q}$ used by the QA-Checker is defined as:
$\mathcal{Q}(Q_{i},A_{i})$ $=\alpha\cdot\text{Relevance}(Q_{i},A_{i})$
$\quad+\beta\cdot\text{Specificity}(A_{i})$
$\quad+\gamma\cdot\text{Coherence}(A_{i})$
where:
  • $Q_{i}$ and $A_{i}$ represent the question and answer at the $i$-th iteration of the conversation.
  • $\text{Relevance}(Q_{i},A_{i})$ measures how well the answer $A_{i}$ addresses the key points and intent of the question $Q_{i}$, computed as:
    $\text{Relevance}(Q_{i},A_{i})=\frac{\vec{Q_{i}}\cdot\vec{A_{i}}}{|\vec{Q_{i}}||\vec{A_{i}}|}$
    where $\vec{Q_{i}}$ and $\vec{A_{i}}$ are vector representations of $Q_{i}$ and $A_{i}$.
  • $\text{Specificity}(A_{i})$ assesses how specific and detailed the answer $A_{i}$ is, calculated as:
    $A_{i}=\frac{\sum_{t\in\text{ContentWords}(A_{i})}\text{TechnicalityScore}(t)}{\text{Length}(A_{i})}$
[…]
  • $\text{Coherence}(A_{i})$ evaluates the logical flow and structural coherence of the answer $A_{i}$, computed as:
    $\text{Coherence}(A_{i})=$ $\alpha\cdot\text{DiscourseConnectives}(A_{i})$
    $+\beta\cdot\text{CoreferenceConsistency}(A_{i})$
    $+\gamma\cdot\text{AnswerPatternAdherence}(A_{i})$
[…]
$\alpha$, $\beta$, and $\gamma$ are non-negative weights that sum to 1, with $\alpha=\beta=\gamma$.
Appendix B Complete Related Work
Automating Code Review Activities Our focus included detecting source code vulnerabilities, ensuring style alignment, and maintaining commit message and code consistency. Other studies explore various aspects of code review. Hellendoorn et al. Hellendoorn et al. (2021) addressed the challenge of anticipating code change positions. Siow et al. Siow et al. (2020) introduced CORE, employing multi-level embeddings for code modification semantics and retrieval-based review suggestions. Hong et al. Hong et al. (2022) proposed COMMENTFINDER, a retrieval-based method for suggesting comments during code reviews. Tufano et al. Tufano et al. (2021) designed T5CR with SentencePiece, enabling work with raw source code without abstraction. Li et al. Li et al. (2022) developed CodeReviewer, focusing on code diff quality, review comment generation, and code refinement using the T5 model. Recently, large language models have been incorporated; Lu et al. Lu et al. (2023) fine-tuned LLama with prefix tuning for LLaMA-Reviewer, using parameter-efficient fine-tuning and instruction tuning in a code-centric domain.
Collaborative AI Collaborative AI refers to artificial intelligent systems designed to achieve shared goals with humans or other AI systems. Previous research extensively explores the use of multiple LLMs in collaborative settings, as demonstrated by Talebirad et al. Talebirad and Nadiri (2023) and Qian et al. Qian et al. (2023). These approaches rely on the idea that inter-agent interactions enable LLMs to collectively enhance their capabilities, leading to improved overall performance. The research covers various aspects of multi-agent scenarios, including collective thinking, conversation dataset curation, sociological phenomenon exploration, and collaboration for efficiency. Collective thinking aims to boost problem-solving abilities by orchestrating discussions among multiple agents. Researchers like Wei et al. Wei et al. (2023) and Li et al. Li et al. (2023a) have created conversational datasets through role-playing methodologies. Sociological phenomenon investigations, such as Park et al. Park et al. (2023)’s work, involve creating virtual communities with rudimentary language interactions and limited cooperative endeavors. In contrast, Akata et al. Akata et al. (2023) scrutinized LLM cooperation through orchestrated repeated games. Collaboration for efficiency, proposed by Cai et al. Cai et al. (2023), introduces a model for cost reduction through large models as tool-makers and small models as tool-users. Zhang et al. Zhang et al. (2023) established a framework for verbal communication and collaboration, enhancing overall efficiency. However, Li et al. Li et al. (2023a) and Qian et al. Qian et al. (2023), presenting a multi-agent framework for software development, primarily relied on natural language conversations, not standardized software engineering documentation, and lacked advanced human process management expertise. Challenges in multi-agent cooperation include maintaining coherence, avoiding unproductive loops, and fostering beneficial interactions. Our approach emphasizes integrating advanced human processes, like code review in software maintenance, within multi-agent systems.
Appendix C Experimental Details
In our work, the maximum number of conversation rounds is set as 10.
C.1 Role Definition
Six roles are defined as shown in Figure 5.
Figure 5: Specialization of six main characters in CodeAgent.
Apart from that, for the QA-checker in CodeAgent, we define an initial prompt for it, which is shown as follows:
[Uncaptioned image] I’m the QA-Checker, an AI-driven agent specializing in ensuring quality and coherence in conversational dynamics, particularly in code review discussions at CodeAgent. My primary role involves analyzing and aligning conversations to maintain topic relevance, ensuring that all discussions about code commits and reviews stay focused and on track. As a sophisticated component of the AI system, I apply advanced algorithms, including Chain-of-Thought reasoning and optimization techniques, to evaluate and guide conversational flow. I am adept at identifying and correcting topic drifts, ensuring that every conversation adheres to its intended purpose. My capabilities extend to facilitating clear and effective communication between team members, making me an essential asset in streamlining code review processes and enhancing overall team collaboration and decision-making.
C.2 Execute Time Across Languages
As depicted in the data, we observe a significant trend in the average execution time for code reviews in CodeAgent across various programming languages. The analysis includes nine languages: Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. For each language, the average execution time of code reviews for both merged and closed pull requests (PRs) is measured. The results, presented in Figure 6, indicate that, on average, the execution time for merged PRs is longer than that for closed PRs by approximately 44.92 seconds. This considerable time difference can be attributed to several potential reasons. One primary explanation is that merged PRs likely undergo a more rigorous and detailed review process. They are intended to be integrated into the main codebase, and as such, contributors might be requested to update their commits in the PRs more frequently to adhere to the project’s high-quality standards. On the other hand, closed PRs, which are not meant for merging, might not require such extensive review processes, leading to shorter review times on average, which may also be the reason they are not merged into main projects.
Figure 6: Execution time with CodeAgent across different language (count unit: second).
Appendix D Comparative Analysis of QA-Checker AI System and Recursive Self-Improvement Systems
In this section, we will delve into the differences between QA-Checker and self-improvement systems Hong et al. (2023), and underscore the importance of the QA-Checker in role conversations.
D.1 Comparison Table
We begin with a comparative overview presented in Table 6.
Table 6: Comparative Overview of QA-Checker AI System and Recursive Self-Improvement Systems
Feature/System QA-Checker AI System Recursive Self-Improvement System
Application Focus
Specialized for QA tasks with
precise task execution
Broad scope, covering various dimensions like
software development and learning algorithms
Learning Mechanism
Advanced optimization techniques
for iterative improvement in QA
Multi-level learning: learning, meta-learning,
and recursive self-improvement
Scope of Improvement
Focused on individual capability
in specific QA tasks
Enhances the entire system, including multi-agent
interactions and communication protocols
Experience Integration
Based on mathematical models
to optimize answer quality
Utilizes experiences from past projects to improve
overall performance
D.2 Differences and Implications
The key differences between these systems lie in their application scope, learning mechanisms, and improvement scopes. The QA-Checker is highly specialized, focusing on QA tasks with efficiency and precision. In contrast, recursive self-improvement systems boast a broader application range and adaptability, integrating experiences from diverse projects for systemic improvements.
D.3 Importance of QA-Checker in Role Conversations
In the context of role conversations, the QA-Checker plays a pivotal role. Its specialized nature makes it exceptionally adept at handling specific conversational aspects, such as accuracy, relevance, and clarity in responses. This specialization is crucial in domains where the quality of information is paramount, ensuring that responses are not only correct but also contextually appropriate and informative.
Furthermore, the efficiency of the QA-Checker in refining responses based on advanced optimization techniques makes it an invaluable tool in dynamic conversational environments. It can quickly adapt to the nuances of a conversation, providing high-quality responses that are aligned with the evolving nature of dialogue.
D.4 Conclusion
While recursive self-improvement systems offer broad adaptability and systemic learning, the QA-Checker stands out in its specialized role in QA tasks, particularly in role conversations. Its focused approach to improving answer quality and its efficiency in handling conversational nuances make it an essential component in AI-driven communication systems.
Appendix E Capabilities Analysis between CodeAgent and Other Methods
Compared to open-source baseline methods such as AutoGPT and autonomous agents such as ChatDev and MetaGPT, CodeAgent offers functions for code review tasks: consistency analysis, vulnerability analysis, and format analysis. As shown in Table 7, our CodeAgent encompasses a wide range of abilities to handle complex code review tasks efficiently. Incorporating the QA-Checker self-improved module can significantly improve the conversation generation between agents and contribute to the improvement of code review. Compared to COT, the difference and the advantages of CodeAgent with QA-Checker are shown in Section D.
Table 7: Comparison of capabilities for CodeAgent and other approaches. ‘✓’ indicates the presence of a specific feature in the corresponding framework, ‘✗ is absence. ChatDev and MetaGPT are two representative multi-agent frameworks, GPT is a kind of single-agent framework, and CodeBert is a representative pre-trained model.
Approaches Consistency Analysis Vulnerability Analysis Format Analysis Code Revision COT QA-Checker
ChatDev Qian et al. (2023)
MetaGPT Hong et al. (2023)
GPT OPENAI (2022)
CodeBert Feng et al. (2020)
CodeAgent
Appendix F Dataset
Previous Dataset
[…]
New Dataset Design and Collection
To enhance our model evaluation and avoid data leakage, we curated a new dataset, exclusively collecting data from repositories created after April 2023. This approach ensures the evaluation of our CodeAgent model on contemporary and relevant data, free from historical biases. The new dataset is extensive, covering a broad spectrum of software projects across nine programming languages.
(a) Positive and negative data of both merged and closed commits across 9 languages on CA task (Sec 2.1).
(b) Positive and negative data of both merged and closed commits across 9 languages on FA task (Sec 2.1).
Figure 7: Distribution of positive, negative of both merged and closed data across 9 languages, including ‘python’, ‘java’, ‘go’, ‘c++’, ‘javascript’, ‘c’, ‘c#’, ‘php’, ‘ruby’.
Dataset Description
Figure 8: Comparative Visualization of Merged and Closed Commit Counts Across Various Programming Languages
Our dataset, illustrated in Fig. 8, encapsulates a detailed analysis of consistency and format detection in software development, spanning various programming languages. It includes CA (consistency between commit and commit message (See Sec 2.1)) and FA (format consistency between commit and original (See Sec 2.1)) data, segmented into positive and negative samples based on the merged and closed status of pull requests. For example, in Python, the dataset comprises 254 merged and 35 closed negative CA samples, alongside 803 merged and 213 closed positive CA samples, with corresponding distributions for other languages like Java, Go, C++, and more. Similarly, the FA data follows this pattern of positive and negative samples across languages. Figure 7 graphically represents this data, highlighting the distribution and comparison of merged versus closed samples in both CA and FA categories for each language. This comprehensive dataset, covering over 3,545 commits and nearly 2,933 pull requests from more than 180 projects, was meticulously compiled using a custom crawler designed for GitHub API interactions, targeting post-April 2023 repositories to ensure up-to-date and diverse data for an in-depth analysis of current software development trends.
Table 8: Statistics of Studied Datasets.
Dataset Statistics #Train #Valid #Test
Trans-Review 13,756 1,719 1,719
AutoTransform 118,039 14,750 14,750
T5-Review 134,239 16,780 16,780
Appendix G Key Factors Leading to Vulnerabilities
The following table outlines various key factors that can lead to vulnerabilities in software systems, along with their descriptions. These factors should be carefully considered and addressed to enhance the security of the system.
No. Vulnerability Factor Description
1 Insufficient Input Validation Check for vulnerabilities like SQL injection, Cross-Site Scripting (XSS), and command injection in new or modified code, especially where user input is processed.
2 Buffer Overflows Particularly in lower-level languages, ensure that memory management is handled securely to prevent overflows.
3 Authentication and Authorization Flaws Evaluate any changes in authentication and authorization logic for potential weaknesses that could allow unauthorized access or privilege escalation.
4 Sensitive Data Exposure Assess handling and storage of sensitive information like passwords, private keys, or personal data to prevent exposure.
5 Improper Error and Exception Handling Ensure that errors and exceptions are handled appropriately without revealing sensitive information or causing service disruption.
6 Vulnerabilities in Dependency Libraries or Components Review updates or changes in third-party libraries or components for known vulnerabilities.
7 Cross-Site Request Forgery (CSRF) Verify that adequate protection mechanisms are in place against CSRF attacks.
8 Unsafe Use of APIs Check for the use of insecure encryption algorithms or other risky API practices.
9 Code Injection Look for vulnerabilities related to dynamic code execution.
10 Configuration Errors Ensure that no insecure configurations or settings like open debug ports or default passwords have been introduced.
11 Race Conditions Analyze for potential data corruption or security issues arising from race conditions.
12 Memory Leaks Identify any changes that could potentially lead to memory leaks and resource exhaustion.
13 Improper Resource Management Check resource management, such as proper closure of file handles or database connections.
14 Inadequate Security Configurations Assess for any insecure default settings or unencrypted communications.
15 Path Traversal and File Inclusion Vulnerabilities Examine for risks that could allow unauthorized file access or execution.
16 Unsafe Deserialization Look for issues that could allow the execution of malicious code or tampering with application logic.
17 XML External Entity (XXE) Attacks Check if XML processing is secure against XXE attacks.
18 Inconsistent Error Handling Review error messages to ensure they do not leak sensitive system details.
19 Server-Side Request Forgery (SSRF) Analyze for vulnerabilities that could be exploited to attack internal systems.
20 Unsafe Redirects and Forwards Check for vulnerabilities leading to phishing or redirection attacks.
21 Use of Deprecated or Unsafe Functions and Commands Identify usage of any such functions and commands in the code.
22 Code Leakages and Hardcoded Sensitive Information Look for hardcoded passwords, keys, or other sensitive data in the code.
23 Unencrypted Communications Verify that data transmissions are securely encrypted to prevent interception and tampering.
24 Mobile Code Security Issues For mobile applications, ensure proper handling of permission requests and secure data storage.
25 Cloud Service Configuration Errors Review any cloud-based configurations for potential data leaks or unauthorized access.
Appendix H Data Leakage Statement
As the new dataset introduced in Section F, the time of the collected dataset is after April 2023, avoiding data leakage while we evaluate CodeAgent on codeData dataset.
Appendix I Algorithmic Description of CodeAgent Pipeline with QA-Checker
This algorithm demonstrates the integration of QA-Checker within the CodeAgent pipeline, employing mathematical equations to describe the QA-Checker’s iterative refinement process.
Algorithm 1 Integrated Workflow of CodeAgent with QA-Checker
Input: Code submission, commit message, original files
Output: Refined code review document
Initialize phase $p=1$
while $p\leq 4$ do
     Switch: Phase $p$
     Case 1: Basic Info Sync
     Conduct initial information analysis
     Update: $p=2$
     Case 2: Code Review
     Perform code review with Coder and Reviewer
     Update: $p=3$
     Case 3: Code Alignment
     Apply code revisions based on feedback
     Update: $p=4$
     Case 4: Document
     Finalize review document
     Update: $p=5$ (End)
     QA-Checker Refinement (Applies in Cases 2 and 3)
     Let $Q_{i}$ be the current question and $A_{i}$ the current answer
     Evaluate response quality: $qScore=\mathcal{Q}(Q_{i},A_{i})$
     if $qScore$ below threshold then
         Generate additional instruction $aai$
         Update question: $Q_{i+1}=Q_{i}+aai$
         Request new response: $A_{i+1}$
     end if
end while
Return: Refined code review document
In this algorithm, $\mathcal{Q}(Q_{i},A_{i})$ represents the quality assessment function of the QA-Checker, which evaluates the relevance and accuracy of the answer $A_{i}$ to the question $Q_{i}$. If the quality score $qScore$ is below a predefined threshold, the QA-Checker intervenes by generating an additional instruction $aai$ to refine the question, prompting a more accurate response in the next iteration.
Appendix J Detailed Performance of CodeAgent in Various Languages on VA task
In our comprehensive analysis using CodeAgent, as detailed in Table 9, we observe a diverse landscape of confirmed vulnerabilities across different programming languages. The table categorizes these vulnerabilities into ‘merged’ and ‘closed’ statuses for languages such as Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. A significant finding is a markedly high number of ‘merged’ vulnerabilities in Python, potentially reflective of its extensive application or intrinsic complexities leading to security gaps. Conversely, languages like Go, Ruby, and C exhibit notably lower counts in both categories, perhaps indicating lesser engagement in complex applications or more robust security protocols. Table 9 that the ‘closed’ category consistently presents lower vulnerabilities than ‘merged’ across most languages, signifying effective resolution mechanisms. However, an exception is noted in C, where ‘closed’ counts surpass those of ‘merged’, possibly indicating either delayed vulnerability identification or efficient mitigation strategies. Remarkably, the Rateclose is generally observed to be higher than Ratemerge across the languages, exemplifying a significant reduction in vulnerabilities post-resolution. For example, Python demonstrates a Ratemerge of 14.00% against a higher Rateclose of 18.16%. This trend is consistent in most languages, emphasizing the importance of proactive vulnerability management. The Rateavg, representing the proportion of confirmed vulnerabilities against the total of both merged and closed items, further elucidates this point, with C++ showing the highest Rateavg at 16.49%. These insights not only underline the diverse vulnerability landscape across programming languages but also highlight the adeptness of CodeAgent in pinpointing and verifying vulnerabilities in these varied contexts.
Table 9: Vulnerable problems (#) found by CodeAgent. Ratemerge means the value of confirmed divided by the total number in the merged and Rateclose is the value of confirmed divided by the total number in the closed. Rateavg is the value of the confirmed number divided by the total number of the merged and closed.
CodeAgent Python Java Go C++ JavaScript C C# PHP Ruby
merged (total#) 1,057 287 133 138 280 114 206 173 202
merged (confirmed#) 148 17 11 19 34 9 21 28 20
Ratemerge 14.00% 5.92% 8.27% 13.77% 12.14% 7.89% 10.19% 16.18% 9.90%
closed (total#) 248 97 74 56 112 146 62 105 55
closed (confirmed#) 45 10 5 13 16 26 7 15 5
Rateclose 18.16% 10.31% 6.76% 23.2% 14.29% 17.81% 11.29% 14.29% 9.09%
Total number (#) 1,305 384 207 194 392 260 268 278 257
Total confirmed (#) 193 27 16 32 50 35 28 43 25
Rateavg 14.79% 7.03% 7.73% 16.49% 12.76% 13.46% 10.45% 14.47% 9.73%
Appendix K More detailed experimental results on CA and FA tasks
Detailed experimental results of CA are shown in Figure 9 and Figure 10. Detailed experimental results of FA are shown in Figure 11 and Figure 12.
Figure 9: Comparison of models on the merged data across 9 languages on CA task.
Figure 10: Comparison of models on the closed data across 9 languages on CA task.
Figure 11: Comparison of models on the merged data across 9 languages on FA task.
Figure 12: Comparison of models on the closed data across 9 languages on FA task.
Appendix L Case Study
As shown in Table 10, we can easily localize the figure numbers of case studies for specific programming languages.
L.1 Performance on 9 languages
Table 10: Correlation Table between specific programming language and case study.
Programming
Language
Figure No.
Python 13
Java 14
Go 15
C++ 16
JavaScript 17
C 18
C# 19
php 20
Ruby 21
Figure 13: Example in Python project with sha value: “0d3b94bbd7528ed8649cdab4370a62e508bad82c"
Figure 14: Example in Java project with sha value: “6e3c6d17d943f5ac70b421653eb167e0c34b119f"
Figure 15: Example in Go project with sha value: “a6b00b097b1a8140bd8b60177247dc830374c8ed"
Figure 16: Example in C++ project with sha value: “09e88c7c20347cdf513323fa8c616aba0a1b9ea7"
Figure 17: Example in JavaScript project with sha value: “6e9a80d93b470aef4978677ed0b7c9996fddbe20"
Figure 18: Example in C project with sha value: “5f93cf43dd6f809d7927abb79884285ad77e8a58"
Figure 19: Example in C# project with sha value: “0e231c7a81b318e9eade972f7b877e66128ed67d"
Figure 20: Example in PHP project with sha value: “6679f059b9a0972a90df198471188da652f5c547"
Figure 21: Example in Ruby project with sha value: “584f72e7f4c65066ccbd2183af76bf380b6eb974"
L.2 Difference of CodeAgent-3.5 and CodeAgent-4.0
CodeAgent-3.5 and CodeAgent-4.0 in this paper has no difference in general code review, however, CodeAgent-4.0 is more powerful in processing long input sequences and logic reasoning. As shown in Figure 22, we take one example of consistency detection between commit and commit message and find that CodeAgent-4.0 diffs from CodeAgent-3.5 in the detailed explanation. CodeAgent-3.5 output a report with 15k lines while CodeAgent-4.0 outputs a report with more than 17.7k lines. Detailed data is shown in https://zenodo.org/records/10607925.
Figure 22: Comparison of code review between CodeAgent-3.5 and CodeAgent-4.0 on one Java project with sha value: “6e3c6d17d943f5ac70b421653eb167e0c34b119f"
Appendix M Ablation study
In this section, we evaluate the performance of different parts in CodeAgent in vulnerability analysis. CodeAgent is based on chain-of-thought (COT) and large language model (a.k.a. GPT). As shown in Section 4.1, CodeAgent outperforms baselines (a.k.a. CodeBERT, GPT-3.5, GPT-4.0) across 9 different languages. The performance mainly comes from the combination of COT and QA-Checker. Thus, we design an additional version called CodeAgent w/o, which means CodeAgent without QA-Checker. Then, we use CodeAgent w/o to do vulnerability analysis and compare with CodeAgent. We first discuss about the result of CodeAgent w/o and then discuss about comparison between CodeAgent and CodeAgent w/o.
Overview of Vulnerabilities in CodeAgent w/o
Table 11 presents the findings of CodeAgent w/o, a variant of the original CodeAgent, in identifying vulnerabilities across different programming languages. The table showcases the number of ‘merged’ and ‘closed’ vulnerabilities in languages such as Python, Java, Go, C++, JavaScript, C, C#, PHP, and Ruby. Notably, Python leads in the ‘merged’ category with a total of 1,057 cases, of which 140 are confirmed, yielding a Ratemerge of 13.25%. In contrast, languages like Go and Ruby show lower vulnerability counts in both ‘merged’ and ‘closed’ categories. The table also includes Rateclose and Rateavg, providing insights into the effectiveness of vulnerability management across these languages.
Detailed Comparison between CodeAgent and CodeAgent w/o
Comparing the findings in Table 11 with those in Table 9, we observe some notable differences in vulnerability detection by CodeAgent and CodeAgent w/o. While the overall trend of higher ‘merged’ vulnerabilities in Python and lower counts in Go and Ruby remains consistent, Table 11 shows a slight reduction in the Ratemerge for most languages, suggesting a more conservative confirmation approach in CodeAgent w/o. Similarly, Rateclose and Rateavg values in Table 11 generally indicate a lower proportion of confirmed vulnerabilities compared to Table 9, reflecting potentially different criteria or efficacy in vulnerability assessment. These variations highlight the impact of QA-Checker in CodeAgent.
Table 11: Vulnerable problems (#) found by CodeAgent w/o
CodeAgent Python Java Go C++ JavaScript C C# PHP Ruby
merged (total#) 1,057 287 133 138 280 114 206 173 202
merged (confirmed#) 140 17 10 12 28 9 21 28 17
Ratemerge 13.25% 5.92% 7.52% 8.70% 10.00% 7.89% 10.19% 16.18% 8.42%
closed (total#) 248 97 74 56 112 146 62 105 55
closed (confirmed#) 36 9 5 12 16 26 7 15 5
Rateclose 14.52% 9.28% 6.76% 21.43% 14.29% 17.81% 11.29% 14.29% 9.09%
Total number (#) 1,305 384 207 194 392 260 268 278 257
Total confirmed (#) 176 26 15 24 44 35 28 43 22
Rateavg 13.49% 6.77% 7.25% 12.37% 11.22% 13.46% 10.45% 15.47% 8.56%
Appendix N Cost statement
As shown in Table 12, CodeAgent-4 has a higher query time and cost compared to CodeAgent-3.5 due to its increased complexity and capabilities. We acknowledge that the integration of AI models and multi-agent systems may introduce complexity and require specialized knowledge for implementation.
Table 12: Summarizes the average query time and cost per code review for CodeAgent-3.5 and CodeAgent-4.
Model Query Time(min) Cost in USD
CodeAgent-3.5 3 0.017
CodeAgent-4 5 0.122
Appendix O Tool
We develop a website for CodeAgent, which is shown in Figure 23, and it is also accessable by visiting following link:
Figure 23: website of CodeAgent